### v2

# Verifiably Following Complex Robot Instructions with Foundation Models

Abstract
Enabling mobile robots to follow complex natural language instructions is an important yet challenging problem. People want to flexibly express constraints, refer to arbitrary landmarks and verify behavior when instructing robots. Conversely, robots must disambiguate human instructions into specifications and ground instruction referents in the real world. We propose Language Instruction grounding for Motion Planning (LIMP), an approach that enables robots to verifiably follow expressive and complex open-ended instructions in real-world environments without prebuilt semantic maps. LIMP constructs a symbolic instruction representation that reveals the robot’s alignment with an instructor’s intended motives and affords the synthesis of robot behaviors that are correct-by-construction. We perform a large scale evaluation and demonstrate our approach on 150 instructions in five real-world environments showing the generality of our approach and the ease of deployment in novel unstructured domains. In our experiments, LIMP performs comparably with state-of-the-art LLM task planners and LLM code-writing planners on standard open vocabulary tasks and additionally achieves 79% success rate on complex spatiotemporal instructions while LLM and Code-writing planners both achieve 38%. See supplementary materials and demo videos at robotlimp.github.io.
Keywords: Foundation Models, Mobile Manipulation, Task and Motion Planning
1 Introduction
Robots need a rich understanding of natural language to be usable by non-experts in unstructured environments. People, on the other hand, need to be able to verify that a robot has understood a given instruction and will act appropriately. However, natural language instructions can be quite complex, often featuring ambiguous phrasing, intricate spatiotemporal constraints, and unique referents. Consider the instruction: “Go to the building with windows while avoiding the red picnic tables, then bring me the book between the pine tree and the graffiti wall”, to solve this task a robot must ground open-vocabulary referents (“building with windows”, “red picnic tables”, “book”, “pine tree”, “graffiti wall”), adhere to temporal constraints (“then bring”, “while avoiding”), and disambiguate multiple instances of an object category using descriptive spatial information (“book between the pine tree and the graffiti wall”). The rise of foundation models in robotics [1, 2] offers a path to achieving such long-horizon goals, and indeed some approaches for robot instruction following have been proposed, but have largely been focused on navigation [3, 4, 5, 6, 7]. These methods, broadly classified under object goal navigation [8], enable navigation to instances of an object category, but are limited in localizing spatial references and disambiguating object instances based on descriptive language. Other works [9, 10, 11] that extend instruction following to mobile manipulation have been limited to tasks with simple temporal constraints expressed in unambiguous language. Existing efforts typically rely on Large Language Models (LLMs) as complete planners, bypassing intermediate symbolic representations that may provide verification of correctness prior to task execution. Alternative approaches that leverage code-writing LLMs [5, 6, 12] are susceptible to errors in generated code which may lead to unsafe/unaligned robot behaviors. Borrowing ideas from formal methods to map between natural language and specification languages like temporal logics [13] provides a robust framework for language disambiguation, handling complex temporal constraints, and behavior verification. However, prior works along this line require prebuilt semantic maps with discrete sets of prespecified referents/landmarks from which instructions can be constructed [7, 14, 15].
We propose Language Instruction grounding for Motion Planning (LIMP), a method that leverages foundation models and temporal logics to dynamically generate instruction-conditioned semantic maps that enable robots to construct verifiable controllers for following navigation and mobile manipulation instructions with open vocabulary referents and complex spatiotemporal constraints. In a novel environment, LIMP first constructs a 3D representation of the environment via SLAM, then leverages large language models to translate complex natural language instructions into temporal logic specifications with a novel composable syntax that enables referent disambiguation. Instruction referents are then detected and grounded using Visual Language Models (VLMs) and spatial reasoning. Finally, a task and motion plan is synthesized to progressively guide the robot through the subgoals required to satisfy the instruction, as shown in Figure LABEL:fig:splash. In summary, we make the following contributions: (1) A modular framework for translating expressive natural language instructions into temporal logic, grounding instruction referents, and executing the command via task and motion planning (TAMP). (2) A spatial grounding method for detecting and localizing open vocabulary objects with spatial constraints in 3D metric maps. (3) A TAMP algorithm that localizes regions of interest (goal/avoidance zones) and synthesizes constraint-satisfying motion plans to progressively solve long-horizon tasks.
2 Related Works
In this section, we briefly highlight the most relevant comparable works in the space of visual scene understanding, instruction following from natural language, and task and motion planning. We offer a more detailed literature review in our appendix A2.
Visual Scene Understanding: NLMap [10] presents a scene representation that grounds open-vocabulary language queries to spatial locations via pre-trained VLMs. While NLMap is suitable for handling complex descriptions of individual objects, it is fundamentally unable to handle instructions involving complex constraints between multiple objects since it does not capture object relationships. LIMP handles these more complicated instructions by using a novel spatial grounding module to incorporate spatial relationships between objects, and leverages task and motion planners to satisfy these constraints.
Natural Language Instruction Following: Lang2LTL [16] is a multi-stage LLM-based approach that performs entity extraction and replacement to translate natural language instructions into a temporal logic expression. However, Lang2LTL relies on a prebuilt semantic map that grounds expression symbols, whereas our approach uses a novel scene representation that supports open-vocabulary language and generates the relevant landmarks based on the open-vocabulary instruction.
Task and Motion Planning: Action-Oriented Semantic Maps (AOSMs) [17] are a class of semantic maps that include additional models in space for where the robot can execute manipulation skills (represented as symbols), and are designed to work in conjunction with TAMP solvers for mobile manipulation domains. Our scene representation is similar to an AOSM since it captures spatial information about semantic objects of interest, and is compatible with TAMP solvers. However, once an AOSM is generated for a scene, there is only a closed-set of goals that can be planned for, whereas our approach can handle open-vocabulary task specifications.
3 Background
LIMP translates natural language instructions into temporal logic specifications for verifiable task and motion planning. Our approach is compatible with various specification languages and planning frameworks, however we choose Linear Temporal Logic (LTL) [18] to represent instructions due to its proven expressivity in representing a wide range of complex robot mission requirements [19]. LTL offers a rich and expressive syntax for defining the temporal properties of systems. Specifications consist of atomic propositions that reflect an aspect of a system’s state and the logical operators: negation ($\neg$), conjunction ($\land$), disjunction ($\lor$), implication ($\rightarrow$), in addition to temporal operators: next ($\mathcal{X}$), until ($\mathcal{U}$), globally ($\mathcal{G}$) and finally ($\mathcal{F}$). Despite its expressivity, LTL has seen limited use due to the expert knowledge required to construct specifications, however recent works have seen significant success removing this barrier by directly translating natural language into LTL [7, 20, 14, 21, 22, 23].
Behavior Verification: Expressing instructions as temporal logic specifications allows us to verify the correctness of generated plans a priori. However, instead of secondary explicit verification via approaches such as model checking, we follow conventional wisdom [24] and directly use specifications to synthesize plans that are correct-by-construction [25, 26].
4 Problem Definition
Given a natural language instruction $l$, our goal is to synthesize and sequence navigation and manipulation behaviors to produce a policy that satisfies the temporal and spatial constraints expressed in $l$. Spatial constraints determine the success of the task based on the sequence of poses the robot traversed during task execution, and temporal constraints determine the sequencing of spatial constraints as a function of task progression. We assume that a robot equipped with an RGBD camera has navigated around some space and captured a collection of images along with camera poses, with this data we build a metric map $m$ (e.g: pointcloud, 3D voxelgrid) of the environment. This map defines the space of possible $SE(3)$ poses $P$ and can be used to localize the robot (i.e: estimate $p_{robot}\in P$). Noteably, unlike previous work that leverage temporal logics [16], we do not assume access to a semantic map with the locations of objects or predicates prespecified. Instead, we leverage two types of foundation models: a task-agnostic visual language model $\sigma$ that given an image and text provides a bounding box/segmentation of the image based on the text, and an auto-regressive large language model $\psi$ that samples likely language tokens based on a history of tokens.
Navigation: We formalize navigation as object-goal oriented continuous path planning problems, where the goal is to generate paths to a goal pose set $P_{goals}\subset P$ while staying in feasible regions ($P_{feasible}\subset P$) and avoiding infeasible regions ($P_{infeasible}=P_{feasible}^{C}$). Infeasible regions include environment obstacles as well as dynamically determined semantic regions that violate constraints in the instruction $l$.
Manipulation: We formalize interactive manipulation behaviours as options [27] parameterized by objects. Consider an object parameter $\theta$ that parameterizes an option $O_{\theta}=(I_{\theta},\pi_{\theta},\beta_{\theta})$, the initiation set, policy, and termination condition are functions of both the robot pose $P$ and $\theta$. The initiation set $I_{\theta}$ denotes the global reference frame robot positions and object-centric attributes––such as object size––that determine if the option policy $\pi_{\theta}$ can be executed on the object $\theta$. To execute a manipulation skill on an object, an object-goal navigation behavior must first be synthesized and executed to bring the robot into proximity with the object in question. Our formalism is capable of encoding a wide variety of long-horizon object-centric tasks, but for simplicity, we demonstrate our approach with two types of tasks: multi-object goal navigation and open vocabulary mobile pick-and-place [9, 28]. These tasks require navigation, picking, and placing robot behaviors.
5 Language Instruction Grounding for Motion Planning
Figure 1: [A] LIMP translates natural language instructions into temporal logic expressions, where open-vocabulary referents are applied to predicates that correspond to robot skills––note the context-aware resolution of the phrase “blue one” to the referent “blue_sofa”. [B] Visual language models detect referents, while spatial reasoning disambiguates referent instances to generate a 3D semantic map that localizes instruction-specific referents. [C] Finally, the temporal logic expression is compiled into a finite-state machine, which a bi-level task and motion planner uses with dynamically-generated task progression semantic maps to progressively identify goals and constraints in the environment, and generate a plan that satisfies the high-level task specification.
LIMP interprets expressive natural language instructions and generates instruction-conditioned semantic maps that enable robots to verifiably solve long-horizon tasks with complex spatiotemporal constraints (Figure 1). Our approach has a modular structure consisting of a Language Instruction Module, a Spatial Grounding Module and a Task and Motion Planning Module. The first two modules utilize pretrained foundation models, while the third involves a task and motion planner for sequencing task subgoals and generating constraint-satisfying motion plans.
5.1 Language Instruction Module
Figure 2: An instruction is first translated into a conventional LTL formula $\phi_{l}$ that loosely captures the desired temporal occurrence of referent objects, then into our LTL syntax $\varphi_{l}$ with predicate functions that temporally chain required robot skills parameterized by composible referent descriptors.
In this module, we leverage a large language model $\psi$ to translate a natural language instruction $l$ into a linear temporal logic specification $\varphi_{l}$ with a novel composable syntax that enables referent disambiguation. We achieve this with a two-stage in-context learning strategy, the first stage translates $l$ into a conventional LTL formula $\phi_{l}$ where propositions refer to open-vocabulary objects. The second stage then takes as input $l$ and $\phi_{l}$ to generate a new formula $\varphi_{l}$, with predicate functions that correspond to parameterized robot skills. We define three predicate functions: near, pick and release for the primitive navigation and manipulation skills (picking and placing) required for multi-object goal navigation and mobile pick-and-place. Predicate functions in $\varphi_{l}$ are parameterized by Composable Referent Descriptors (CRDs), our novel propositional expressions that represent specific referent instances by chaining comparators that encode descriptive spatial information. For example the statement “the yellow cabinet above the fridge that is next to the stove”, can be represented with the composible referent descriptor below:
[…]
which specifies that there is a fridge that is next to a stove, and the specific yellow cabinet we desire is above that fridge. CRDs are constructed from a set of 3D spatial comparators [29] defined in our prompting strategy. It is important to note that unlike recent works [9, 30], instructions do not require any specific phrasing syntax nor specific keywords and may feature arbitrary complexity and ambiguity. The LLM $\psi$ directly samples the entire LTL formula $\varphi_{l}$ with predicate functions parameterized by composable referent descriptors with appropriate spatial comparators based on the context of the instruction. Figure 2 presents an example highlighting the result of our two-stage prompting strategy. See appendix section A3 for further details.
LLM Verification: Verifying the LTL formula $\varphi_{l}$ sampled from the LLM is vital as errors in referent extraction and temporal task structure implied by the sampled formula adversely affect accurate instruction following. To that end, the symbol verification node highlighted in Figure 1 leverages the unique properties of LTL to provide high-level human-in-the-loop verification for the extracted instruction referents and temporal task structure. Recent work [31] has addressed providing ISO 61508 [32] standard safety guarantees in robot task execution by translating descriptions of safety constraints from natural language to LTL formulas which are verified to be correct by human experts and used to enforce robot behavior. Similarly, we rely on human checking to verify that the translated formula $\varphi_{l}$ is correct, the symbol verification node implements an interactive dialog system that presents users with the extracted referent CRDs in the LTL formula as well as the implied task structure and reprompts the LLM based on corrective statements of the user to obtain new formulas. Unlike [31] we improve readability and relax the need for experts by directly translating the task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task––back into English statements via a simple deterministic translation scheme. In our experiments (Table 1 & 2), we find that even without this human verification and reprompting step, the initial formulas sampled by our language understanding module impressively encodes the correct referents and implied temporal task structure. See appendix section A3.1 for further details.
5.2 Spatial Grounding Module
Figure 3: [A] Our spatial grounding module leverages a VLM to detect all referent occurrences from prior observations of the environment. [B] We then obtain an initial semantic map with all detected referent instances by backprojecting pixels in segmented referent masks unto the 3D map. [C] We resolve each referent’s spatial comparators with respect to the origin coordinate frame of reference. [D] We finally filter out failing instances to obtain a Referent Semantic Map (RSM) that localizes the exact referent instances described in the instruction.
The spatial grounding module detects and localizes specific instances of objects referenced in a given instruction. To achieve this we extract composable referent descriptors from translated LTL formulae and leverage visual language models [33, 34] to detect and generate segmentation masks of all referent occurrences from the robot’s prior observations of the environment. We then backproject pixels in these masks unto our 3D map to obtain an initial semantic map that localizes all instruction object instances. In the example from Figure 2, all occurrences of green_plush_toy, whiteboard and robot are detected, segmented and backprojected unto the map as shown in Figure 3[a&b].
We filter detected occurrences of referents to obtain the specific object instances described in the given instruction by resolving the 3D spatial comparators captured in each referent’s CRD, recall that CRDs are propositional expressions and can be evaluated as true or false. We define eight (8) spatial comparators (isbetween, isabove, isbelow, isleftof, isrightof, isnextto, isinfrontof, isbehind) to reason about the spatial relationships between referents based on their backprojected 3D positions. Since all backprojected positions are relative to an origin coordinate system, our spatial comparators are resolved from the perspective of this origin position as shown in Figure 3[c]. This type of Relative frame of reference (FoR) when describing spatial relationships between objects, in contrast to an Absolute or Intrinsic FoR, is dominant in English [35], hence a logical choice in our work. For each detected referent, the backprojected 3D position of its center mask pixel is used as its representative position and we resolve a given referent with a spatial description by applying the appropriate spatial comparator to all detected pairs of the desired referent and comparison landmark objects. This filtering process yields a Referent Semantic Map (RSM) that localizes specific object instances described in the original instruction as shown in Figure 3[d]. See appendix section A4 for further details.
VLM Verification: Potential misclassifications from object detector VLMs is the main source of error in this module. We do not address interactively correcting VLM misclassifications as that is out of the scope of this work, however we provide 3D visualizations tools that enable end users to visually inspect and verify that constructed referent semantic maps localized referents correctly.
5.3 Task and Motion Planning Module
Figure 4: [A] A given instruction translated into our LTL syntax $\varphi_{l}$ can be compiled into an equivalent finite-state automaton that captures the temporal constraints of the task. We select a path through this automaton with a strategy that incrementally selects the next progression state from the initial state to the accepting state. We then execute manipulation options and navigation behaviors as required by this high-level task plan. [B] To execute navigation objectives we generate a task progression semantic map (TPSM) that augments the environment with state transition constraints, localizing goal (yellow) and avoidance (red) regions. We convert generated TPSMs into 2D obstacle maps for constraint-aware continuous path planning.
Finally, our TAMP module synthesizes and sequences navigation and manipulation behaviors to produce a plan that satisfies the temporal and spatial constraints expressed in the given instruction.
Progressive Motion Planner: Our TAMP algorithm (Alg.1) first compiles an LTL formula with parameterized robot skills into an equivalent finite-state automaton (Figure 4[a]), which can be used to generate a plan that is guaranteed to be correct-by-construction. We select a path that traverses this automaton from its initial state to its accepting state, this is a high-level task plan that interleaves the navigation and manipulation objectives required for satisfying the instruction. We implement a simple strategy that incrementally selects the next progression state from the initial state until the final accepting state is reached, ensuring the generated plan faithfully obeys all temporal subgoal objectives. As shown in Figure 4[a] each progressive automaton state is connected by transition edges that represent the logical expression that must hold for the transition to occur. Our algorithm performs the required low-level behavior to achieve each transition, for manipulation objective subgoals, we execute the appropriate parameterized skill. For navigation objective subgoals, we dynamically generate Task Progression Semantic Maps (TPSM) that localize goal and constraint regions in the environment and perform continuous path planning using the Fast Marching Tree algorithm (FMT) [36].
Task Progression Semantic Maps: A TPSM (Alg.2) augments a 3D scene with the navigation constraints specified by logical state transition expressions, enabling a robot to perform goal-directed and constraint-aware navigation. Segmented regions of interest captured in a TPSM can be arbitrarily large via a nearness threshold (in meters) that defines what it means to be at or close to an object. This parameter can be specified as a high-level parameter or included in the language instruction module’s prompting strategy, such that an LLM infers what value to use based on the given instruction. Goal regions captured in a generated TPSM along with object-centric constraints, such as object shape and size, form the initiation set for manipulation options. Similar to our spatial grounding module, TPSMs are agnostic to temporal logic task representations and can be used out of the box with a variety of planning approaches to localise regions of interest and enable semantic constraint aware motion planning. We primarily evaluate our approach on a ground mobile robot, hence we transform 3D TPSMs into 2D geometric obstacle maps, where constraint regions are also treated as obstacle regions (Figure 4[b]). Our approach however, is robot agnostic and allows for direct planning in 3D TPSMs for appropriate embodiments, such as drones.
6 Experimental Evaluation
Table 1: Performance comparison of one-shot instruction understanding and spatial resolution.
Approach Referent Resolution Accuracy (Average WER) $\downarrow$ Avoidance Constraint Resolution Accuracy (Average WER) $\downarrow$ Spatial Relationship Resolution Accuracy (Average WER) $\downarrow$
NLMap-Saycan 0.09 0.12 0.05
Code-as-policies 0.22 0.24 0.06
Limp Single Stage Prompting 0.09 0.11 0.05
Limp Two Stage Prompting [Random Embedding] 0.08 0.11 0.03
Limp Two Stage Prompting [Similar Embedding] 0.07 0.04 0.03
Table 2: Performance comparison of one-shot temporal alignment and plan success rate.
Approach Temporal Alignment Accuracy (of instructions) $\uparrow$ Planning Success Rate (of instructions) $\uparrow$
NLMD NLMC OKRB CT CST NLMD NLMC OKRB CT CST
NLMap-Saycan 88% 96% 100% 32% 41% 75% 96% 100% 35% 38%
Code-as-policies 58% 68% 100% 35% 38% 46% 68% 100% 38% 38%
Limp Single Stage Prompting 79% 64% 100% 68% 74% 63% 60% 100% 62% 62%
Limp Two Stage Prompting [Random Embedding] 83% 68% 100% 76% 85% 79% 68% 100% 57% 72%
Limp Two Stage Prompting [Similar Embedding] 88% 80% 100% 76% 92% 79% 76% 100% 65% 79%
Our evaluations test the hypothesis that our system for translating natural language instructions into LTL expressions and dynamically generating semantic maps enables robots to accurately interpret and execute instructions in large-scale environments without prior training in said environments. We focus on three key questions: (1) Can our language module interpret complex ambiguous natural language instructions? (2) Can our spatial grounding module resolve specific object instances described in instructions? (3) Can our TAMP algorithm generate constraint-satisfying plans?
Figure 5: Our interactive reprompting strategy implemented in the symbol verification node regenerates corrective formulas that improve plan success rates with minimal chat turns.
To answer these questions, we compare LIMP with baseline implementations of an LLM task planner (NLMap-Saycan) [10] and an LLM code-writing planner (Code-as-policies) [12], representing state-of-the-art approaches for language-conditioned open ended robot task execution. Both baselines use the same underlying LLM (gpt-4-0613), prompting structure, and in-context learning examples as our language understanding module. For the code-as-policies planner, in-context examples are converted into language model-generated program (LMP) snippets [12]. To ensure competitive performance, we integrate our CRD syntax, spatial grounding module and low-level robot control into these baselines, allowing them to query object positions, use our FMT* path planner, and execute manipulation options. We also evaluate ablations of our two-stage language understanding module, due to the prominence of language understanding on instruction following. In our complete approach, the first stage samples a conventional LTL formula $\phi_{l}$ from an instruction $l$ by dynamically selecting relevant in-context examples from a standard dataset [14] of natural language and LTL pairs based on cosine similarity. Our first ablation selects in-context examples randomly, and our second ablation skips this stage entirely, directly sampling our LTL syntax with parameterized robot skills $\varphi_{l}$.
Figure 6: [A] Sample generated plan for a multi object-goal navigation task. [B] Sample generated plan for a mobile pick-and-place task.
We perform a large-scale evaluation on 150 instructions across five real-world environments. The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [16], and an additional 39 tasks with descriptive spatial constraints and temporal structures (CST). To evaluate instruction understanding, we introduce the following performance metrics: Referent resolution accuracy, Avoidance constraint resolution accuracy, and Spatial relationship resolution accuracy. These metrics use the popular word error rate metric (WER) to measure alignment with the ground-truth. Additionally, we define Temporal alignment accuracy and Planning success rate metrics. A plan is temporally aligned if the sequence of subgoal objectives matches the instructor’s intention, and it is deemed successful if it satisfies all spatial and temporal constraints specified in the instruction. Achieving a high plan success rate requires accurate referent and avoidance constraint resolution, spatial grounding, and temporal alignment. We report the average word error rates for each baseline in Table 1 and the percentage of successful plans and temporally accurate plans in Table 2. In addition to the verification benefits of symbolic planning, LIMP outperforms baselines in most tasksets, especially in complex temporal planning and constraint avoidance. We observe that LLM and code-writing planners generate sequential subgoals effectively, but struggle with temporal constraint adherence. In contrast, our approach ensures each robot step adheres to constraints while achieving subgoals. Additional qualitative results are shown in Figure 6, and robot demonstrations are available on our project website.
7 Limitations and Conclusion
Our study revealed several limitations that merit discussion. LIMP is able to interpret non-finite instructions into LTL formulas however our planner is only able to processes co-safe formulas, which are limited to finite sequences. The accuracy of spatial grounding is dependent on VLM object recognition and depth sensing, with any deficiencies affecting performance. LIMP assumes a consistent environment between mapping and execution, making our approach non-reactive to dynamic changes. Addressing this with editable scene representations is a future goal. Moreover, while our Progressive Motion Planning algorithm is complete, it does not provide any optimality guarantees, future work will focus on enhancing planning efficiency.
Foundation models have the potential to drive the next generation of generally capable robots. Our results suggest leveraging these models for their strengths (LLMs for language, VLMs for vision) in combination with proven approaches for safety, explainability, and verifiable behavior synthesis. This approach for using foundation models in robotics is vital to improve safety and trust in robots that will have to work with and for humans.
Supplementary Materials
Appendix A1 Appendix Summary
These sections presents additional details on our approach for leveraging foundation models and temporal logics to verifiably follow expressive natural language instructions with complex spatiotemporal constraints without prebuilt semantic maps. We encourage readers to visit our website robotlimp.github.io for project summary and demonstration videos.
Appendix A2 Extended Related Works
A2.1 Foundation Models in Robotics
Grounding language referents to entities and actions in the world [37, 38, 39, 40] is challenging in part due to the fact that complex perceptual and behavioral meaning can be constructed from the composition of a wide-range of open-vocabulary components[41, 20, 21, 42, 38]. To address this problem, foundation models have recently garnered interest as an approach for generating perceptual representations that are aligned with language [43, 4, 44, 45, 46, 47]. Because there are an ever-expanding number of ways foundation models are being leveraged for instruction following in robotics (e.g: generating plans [47], code [48], etc.), we focus our review on the most related approaches in two relevant application areas : 1) generating natural language queryable scene representations and 2) generating robot plans for following natural language instruction[16, 49].
Visual scene understanding: The most similar approach for visual scene understanding to ours is NLMap [10], a scene representation that affords grounding open-vocabulary language queries to spatial locations via pre-trained visual language models. Given a sequence of calibrated RGB-D camera images and pre-trained visual-language models, NLMap supports language-queries by 1) segmenting out the most likely objects in the 2D RGB images based on the language queries, and 2) estimating the most likely 3D positions via back-projection of the 2D segmentation masks using the depth data and camera pose. While NLMap is suitable for handling complex descriptions of individual objects (e.g: “green plush toy“), it is fundamentally unable to handle instructions involving complex constraints between multiple objects since it has no way to account for object-object relationships (e.g: “the green plush toy that is between the toy box and door“). LIMP handles these more complicated language instructions by using a novel spatial grounding module to easily incorporate a wide-variety of complex spatial relationships between objects. In addition, our scene representation is compatible with both LLM planners as well as TAMP solvers, whereas NLMap is only compatible with LLM planners.
While NLMap is the most relevant approach to ours, there are other approaches for visual scene understanding and task planning with foundation models which are worth highlighting. VoxPoser [50] leverages the abilities of LLMs to identify affordances and write code for manipulation tasks, along with VLMs complementary abilities to identify open-vocabulary entities in the environment. SayPlan [51] integrates 3D scene graphs with LLM-based planners to bridge the gap between complex, heirarchial scene representations and scalable task planning with open-ended task specifications. Generalizable Feature Fields (GeFF) [52] use an implicit scene representation to support open-world manipulation and navigation via an extension of Neural Radiance Fields (NeRFs) and feature distillation in NeRFs. OK-Robot [30] adopts a system-first approach to solving structured mobile pick-and-place tasks with foundation models by offering an integrated solution to object detection, mapping, navigation and grasp generation. While these methods are related, none of them have all the features of LIMP: 1) Explicit support for both LLM-based planning and off-the-shelf task and motion planning approaches, 2) Verifiable representations for following complex natural language instructions in mobile manipulation domains that involve object-object relationships, and 3) The ability to dynamically generate task-relevant state abstractions (semantic maps) for individual instructions.
Language instruction for robots: Our approach to handling complex natural language instructions involves translating the command into a temporal logic expression. This problem framing allows us to leverage state-of-the-art techniques from machine translation, such as instruction-tuned large language models. Most similar to our approach in this regard is [16], which uses a multi-stage LLM-based approach and finetuning to perform entity-extraction and replacement to translate natural language instructions into temporal logic expressions. However, [16] relies on a prebuilt semantic map that grounds expression symbols, limiting the scope of instructions it can operate since landmarks are predetermined. Instead, our approach interfaces with a novel scene representation that supports open-vocabulary language and generates the relevant landmarks based on the open-vocabulary instruction. Additionally, the symbols in our temporal logic translation correspond to parameterized task relevant robot skills as opposed to propositions of referent entities extracted from instructions.
A2.2 Planning Models in Robotics
Semantic Maps: Semantic maps [53] are a class of scene representations that capture semantic (and typically geometric) information about the world, and can be used in cojunction with planners to generate certain types of complex robot behavior like collision-free navigation with spatial constraints [54, 55]. However, leveraging semantic maps for task planning with mobile manipulators has been challenging since the modeling information needed may highly depend on the robot’s particular skills and embodiment. [17] recently proposed Action-Oriented Semantic Maps (AOSMs), which are a class of semantic maps that include additional models of the regions of space where the robot can execute manipulation skills (represented as symbols). [17] demonstrated that AOSMs can be used as a state representation that supports TAMP solvers in mobile manipulation domains. Our scene representation is similar to an AOSM since it captures spatial information about semantic regions of interest, and is compatible with TAMP solvers, but largely differs in that AOSMs require learning via online interaction with the scene. Instead, our approach leverages foundation models and requires no online learning. Also, once an AOSM is generated for a scene, there is only a closed-set of goals that can be planned for, whereas our approach can handle open-vocabulary task specifications.
Task and Motion Planning: Task and Motion planning approaches are hierarchical planning methods that involve high-level task planning (with a discrete state space) [56] and low-level motion planning (with a continuous state space) [57]. The low-level motion planning problem involves generating paths to goal sets through continuous spaces (e.g: configuration space, cartesian space) with constraints on infeasible regions. When the constraints and dynamics can change, it is referred to as multi-modal motion planning, which naturally induces a high-level planning problem that involves choosing which sequence of modes to plan through, and a low-level planning problem that involves moving through a particular mode. Finding high-level plan skeletons and satisfying low-level assignment values for parameters to achieve goals is a challenging bi-level planning problem[57]. LIMP contains sufficient information to produce a problem and domain description augmented with geometric information for bi-level TAMP solvers like [58, 59].
Appendix A3 Language Instruction Module
We implement a two-stage prompting strategy in our language instruction module to translate natural language instructions into LTL specifications. The first stage translates a given instruction into a conventional LTL formula, where propositions refer to open-vocabulary objects. For any given instruction, we dynamically generate $K$ in-context translation examples from a standard dataset [14] of natural language and LTL pairs, based on cosine similarity with the given instruction. Here is the exact text prompt used:
1You are a LLM that understands operators involved with Linear Temporal Logic (LTL), such as F, G, U, &, |, ~ , etc. Your goal is translate language input to LTL output.
2Input:<generated_example_instruction>
3Output:<generated_example_LTL>
4...
5Input:<given_instruction>
6Output:
Listing 1: Base prompt used to obtain a conventional LTL formula from a natural language query
The second stage takes the given instruction and the LTL response from the first stage as input to generate a new LTL formula with predicate functions that correspond to parameterized robot skills. Skill parameters are instruction referent objects expressed in our novel Composable Referent Descriptor (CRD) syntax. CRDs enable referent disambiguation by chaining comparators that encode descriptive spatial information. We define eight spatial comparators and provide their descriptions as part of the second stage prompt. We find that LLMs conditioned on this information and a few examples are able translate arbitrarily complex instructions with appropriate comparator choices. Here is the exact prompt used:
1You are an LLM for robot planning that understands operators involved with Linear Temporal Logic (LTL), such as F, G, U, &, |, ~ , etc. You have a finite set of robot predicates and spatial predicates, given a language instruction and an LTL formula that represents the given instruction, your goal is to translate the ltl formula into one that uses appropriate composition of robot and spatial predicates in place of propositions with relevant details from original instruction as arguments.
2Robot predicate set (near,pick,release).
3Usage:
4near[referent_1]:returns true if the desired spatial relationship is for robot to be near referent_1.
5pick[referent_1]:can only execute picking skill on referent_1 and return True when near[referent_1].
6release[referent_1,referent_2]:can only execute release skill on referent_1 and return True when near[referent_2].
7Spatial predicate set (isbetween,isabove,isbelow,isleftof,isrightof,isnextto,isinfrontof,isbehind).
8Usage:
9referent_1::isbetween(referent_2,referent_3):returns true if referent_1 is between referent_2 and referent_3.
10referent_1::isabove(referent_2):returns True if referent_1 is above referent_2.
11referent_1::isbelow(referent_2):returns True if referent_1 is below referent_2.
12referent_1::isleftof(referent_2):returns True if referent_1 is left of referent_2.
13referent_1::isrightof(referent_2):returns True if referent_1 is right of referent_2.
14referent_1::isnextto(referent_2):returns True if referent_1 is close to referent_2.
15referent_1::isinfrontof(referent_2):returns True if referent_1 is in front of referent_2.
16referent_1::isbehind(referent_2):returns True if referent_1 is behind referent_2.
17Rules:
18Strictly only use the finite set of robot and spatial predicates!
19Strictly stick to the usage format!
20Compose spatial predicates where necessary!
21You are allowed to modify the structure of Input_ltl for the final Output if it does not match the intended Input_instruction!
22You should strictly only stick to mentioned objects, however you are allowed to propose and include plausible objects if and only if not mentioned in instruction but required based on context of instruction!
23Pay attention to instructions that require performing certain actions multiple times in generating and sequencing the predicates for the final Output formula!
24Example:
25Input_instruction: Go to the orange building but before that pass by the coffee shop, then go to the parking sign.
26Input_ltl: F (coffee_shop & F (orange_building & F parking_sign ) )
27Output: F ( near[coffee_shop] & F ( near[orange_building] & F near[parking_sign] ))
28Input_instruction: Go to the blue sofa then the laptop, after that bring me the brown bag between the television and the kettle on the left of the green seat, I am standing by the sink.
29Input_ltl: F ( blue_sofa & F ( laptop & F ( brown_bag & F ( sink ) ) ) )
30Output: F ( near[blue_sofa] & F ( near[laptop] & F ( near[brown_bag::isbetween(television,kettle::isleftof(green_seat))] & F (pick[brown_bag::isbetween(television,kettle::isleftof(green_seat))] & F ( near[sink] & F ( release[brown_bag,sink] ) ) ) ) ) )
31Input_instruction: Hey need you to pass by chair between the sofa and bag, pick up the bag and go to the orange napkin on the right of the sofa.
32Input_ltl: F ( chair & F ( bag & F ( orange_napkin ) ) )
33Output: F ( near[chair::isbetween(sofa,bag)] & F ( near[bag] & F ( pick[bag] & F ( near[orange_napkin::isrightof(sofa)] ) ) ) )
34Input_instruction: Go to the chair between the green laptop and the yellow box underneath the play toy
35Input_ltl: F ( green_laptop & F ( yellow_box & F ( play_toy & F ( chair ) ) ) )
36Output: F ( near[chair::isbetween(green_laptop,yellow_box::isbelow(play_toy))] )
37Input_instruction: Check the table behind the fridge and bring two beers to the couch one after the other
38Input_ltl: F ( check_table & F ( bring_beer1 ) & F ( bring_beer2 ) & F ( couch ) )
39Output: F ( near[table::isbehind(fridge)] & F ( pick[beer] & F ( near[couch] & F ( release[beer,couch] & F ( near[table::isbehind(fridge)] & F ( pick[beer] & F ( near[couch] & F ( release[beer,couch] ))))))))
40Input_instruction: <given_instruction>
41Input_ltl: <stage1_ltl_response>
42Output:
Listing 2: Second stage prompt to output our LTL syntax with CRD parameterized robot skills
A3.1 Interactive Symbol Verification
Verifying sampled LTL formulas is essential, as such we implement an interactive dialog system that presents users with extracted referent composible referent descriptors (CRDs) in sampled formulas as well as the implied task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task. We translate the task structure into English statements via a simple deterministic strategy that replaces logical connectives and skill predicates from the formula with equivalent English phrases. Users can verify a formula as correct or provide corrective statements which are used to reprompt the LLM to obtain new formulas. Here is the exact text prompt used for reprompting.
1There was a mistake with your output LTL formula: Error with <verification_type>! Consider the clarification feedback and regenerate the correct output for the Input_instruction. Make sure to adhere to all rules and instructions in your original prompt!
2previous_output:<last_response>
3error_clarification: <given_error_clarification>
4correct_output:
Listing 3: Corrective reprompting prompt used to obtain new LTL formulas
As an illustration, the instruction “Bring the green plush toy to the whiteboard in front of it” yields the interactive Referent and Task Structure Verification dialog shown in Figure A.7.
Figure A.7: Sample interactive referent and task structure verification dialog
Appendix A4 Spatial Grounding Module
The spatial grounding module detects and localizes specific instances of objects referenced in a given instruction by first detecting, segmenting and back-projecting all referent occurances and then filtering based on the descriptive spatial details captured by each referent’s composable referent descriptor (CRD). We use the Owl-Vit model [33] to detect bounding boxes of open-vocabulary referents and SAM [34] to generate masks from detected bounding boxes. To illustrate referent filtering via spatial information, consider an example scenario where the goal is to resolve the composable referent descriptor below:
[…]
Let $W=\{w_{1},w_{2},\ldots,w_{n}\}$ and $G=\{g_{1},g_{2},\ldots,g_{m}\}$ represent the set of representative 3D positions of detected whiteboards and green_plush_toys respectively. The cartesian product of these sets enumerates all possible pairs $(w,g)$ for comparison.
$W\times G=\{w,g)\mid w\in W,g\in G\}$ (A.3)
The ‘isinfrontof(w, g)’ comparator is applied to each pair, yielding a subset $S$ that contains only those ‘whiteboards‘ that satisfy the ‘isinfronto’ condition with at least one ‘green_plush_toy’.
$S=\{w\in W\mid\exists g\in G\text{ such that }\text{isinfrontof}(w,g)\text{ istrue}\}$ (A.4)
A4.1 3D Spatial Comparators
Our 3D spatial comparators enable Relative Frame of Reference (FoR) spatial reasoning between referents, based on their backprojected 3D positions. Threshold values in the spatial comparators give users the ability to specify the sensitivity or resolution at which spatial relationships are resolved, we keep all threshold values fixed across all experiments. Below is a description of each spatial comparator.
11. isbetween(referent_1_pos, referent_2_pos, referent_3_pos, threshold): Returns true if referent_1 is within threshold distance from the line segment connecting referent_2 to referent_3, ensuring it lies in the directional path between them without extending beyond.
22. isabove(referent_1_pos, referent_2_pos, threshold): Returns true if the z-coordinate of referent_1 exceeds that of referent_2 by at least threshold’.
33. isbelow(referent_1_pos, referent_2_pos, threshold): Returns true if the z-coordinate of referent_1 is less than that of referent_2 by more than threshold’.
44. isleftof(referent_1_pos, referent_2_pos, threshold): Returns true if the y-coordinate of referent_1 exceeds that of referent_2 by at least threshold’, indicating referent_1 is to the left of referent_2.
55. isrightof(referent_1_pos, referent_2_pos, threshold): Returns true if the y-coordinate of referent_1 is less than that of referent_2 by more than threshold’, indicating referent_1 is to the right of referent_2.
66. isnextto(referent_1_pos, referent_2_pos, threshold): Returns true if the Euclidean distance between referent_1 and referent_2 is less than threshold’, indicating they are next to each other.
77. isinfrontof(referent_1_pos, referent_2_pos, threshold): Returns true if the x-coordinate of referent_1 is less than that of referent_2 by more than threshold’, indicating referent_1 is in front of referent_2.
88. isbehind(referent_1_pos, referent_2_pos, threshold): Returns true if the x-coordinate of referent_1 exceeds that of referent_2 by at least threshold’, indicating referent_1 is behind referent_2.
Listing 4: Implementation description of 3D spatial comparators
Appendix A5 Task and Motion Planning Module
We present pseudocode for our Progressive Motion Planner (Alg.1) and our algorithm for generating Task Progression Semantic Maps (Alg.2). Alg.2 generates a TPSM $\mathcal{M}_{\text{tpsm}}$ by integrating an environment map ($\mathcal{M}$) and a referent semantic map ($\mathcal{M}_{\text{rsm}}$) given a logical transition expression ($\mathcal{T}$), a desired automaton state ($\mathcal{S^{\prime}}$), and a nearness threshold ($\theta$). The algorithm first initializes $\mathcal{M}_{\text{tpsm}}$ with a copy of $\mathcal{M}$ and extracts relevant instruction predicates from $\mathcal{T}$. For each predicate (parameterized skill), the algorithm identifies satisfying referent positions in $\mathcal{M}_{\text{rsm}}$, generates a spherical grid of surrounding points within a radius $\theta$, and assesses how these points affect the progression of the task automaton towards $\mathcal{S^{\prime}}$. These points demarcate regions of interest, and are assigned a value of 1 if they cause the automaton to transition to the desired state, -1 if they lead to a different automaton state or violate the automaton, and 0 if they do not affect the automaton. The points are then integrated into $\mathcal{M}_{\text{tpsm}}$, yielding a semantic map that identifies goal and constraint violating regions.
A5.1 Algorithm Pseudocode
Algorithm 1 Progressive Motion Planning Algorithm
1:procedure PMP($X_{start},\varphi,\mathcal{M},\mathcal{M}_{rsm},\theta$)
2:Input:
3:$X_{start}$: Start position in the environment.
4:$\varphi$: CRD syntax LTL formula specifying task objectives.
$\mathcal{M}$: Environment map.
6:$\mathcal{M}_{rsm}$: Referent semantic map.
7:$\theta$: Nearness threshold.
8:Output:
9:$\Pi$: Generated task and motion plan.
10:     $\mathcal{A}\leftarrow\text{ConstructAutomaton}(\varphi)$
[…]
12:     while $\Pi$.status is active do
13:         while $\mathcal{A}$.state != path.acceptingState do
[…]
17:                  $\Pi.\text{UpdateWithSkill}(\mathcal{T})$
18:                  $\mathcal{A}$.UpdateAutomatonState($\mathcal{S^{\prime}}$)
[…]
20:                  $\mathcal{M}_{params}\leftarrow\mathcal{M},\mathcal{M}_{rsm},\mathcal{T},\mathcal{A},\mathcal{S^{\prime}},\theta$
21:                  $\mathcal{M}_{tpsm}\leftarrow\textsc{GenerateTPSM}(\mathcal{M}_{params})$
22:                  $\mathcal{O}\leftarrow\text{GenerateObstacleMap}(\mathcal{M}_{tpsm})$
23:                  $\text{plan}\leftarrow\text{FMT}^{*}(X_{start},\mathcal{O})$triangleright Path plan
24:                  if plan.exists then
25:                       $\Pi.\text{UpdateWithPlan}(\text{plan})$
[…]
27:                       $\mathcal{A}$.UpdateAutomatonState($\mathcal{S^{\prime}}$)
28:                  else
29:                       $\Pi,\mathcal{A},\text{path}\leftarrow\text{Backtrack}(\Pi,\mathcal{A},\text{path})$
30:                  end if
31:              end if
32:         end while
33:     end while
34:     return $\Pi$
35:end procedure
Algorithm 2 Task Progression Semantic Mapping Algorithm
1:procedure GenerateTPSM($\mathcal{M},\mathcal{M}_{rsm},\mathcal{T},\mathcal{A},\mathcal{S^{\prime}},\theta$)
2:Input:
$\mathcal{M}$: Environment map.
4:$\mathcal{M}_{rsm}$: Referent semantic map.
5:$\mathcal{T}$: Automaton transition expression.
6:$\mathcal{A}$: Task Automaton.
7:$\mathcal{S^{\prime}}$: Desired State.
8:$\theta$: Nearness threshold.
9:Output:
10:$\mathcal{M}_{tpsm}$: Task Progression Semantic Map.
11:     $\mathcal{M}_{tpsm}\leftarrow\text{Copy}(\mathcal{M})$
12:     $\mathcal{P}\leftarrow\text{ExtractRelevantPredicates}(\mathcal{T})$
13:     for $p$ in $\mathcal{P}$ do
14:         $\mathcal{R}\leftarrow\text{QueryPositions}(\mathcal{M}_{rsm},p)$
15:         for $r$ in $\mathcal{R}$ do
16:              $G\leftarrow\left\{g\mid g=r+\delta,\lVert\delta\rVert\leq\theta\right\}$triangleright spherical grid of surrounding points
17:              for $g$ in $G$ do
18:                  $\mathcal{Q}\leftarrow\text{TruePredicatesAt}(g,\mathcal{M}_{rsm},\theta)$
19:                  $\mathcal{S}_{next}\leftarrow\text{ProgressAutomaton}(\mathcal{A},\mathcal{Q})$
20:                  if $\mathcal{S}_{next}=\mathcal{S^{\prime}}$ then
21:                       $g.value\leftarrow 1$triangleright Goal value
22:                  else if $\text{IsUndesired}(\mathcal{S}_{next})$ then
23:                       $g.value\leftarrow-1$triangleright Avoidance value
24:                  else
25:                       $g.\text{value}\leftarrow 0$
26:                  end if
27:              end for
28:              $\text{AddPoints}(\mathcal{M}_{tpsm},G)$
29:         end for
30:     end for
31:     return $\mathcal{M}_{tpsm}$
32:end procedure
Appendix A6 Robot Skills
We define three predicate functions: near, pick and release for the navigation, picking and placing skills required for multi-object goal navigation and mobile pick-and-place. As highlighted in the main paper, we formalize navigation as continuous path planning problems and manipulation as object parameterized options. We discuss navigation at length in the paper, so here we focus on the pick and place manipulation skills.
Pick Skill: Once the robot has executed the near skill and is at the object to be manipulated, we take a picture of the current environment to detect the object using the Owl-Vit model. The robot is guaranteed to be facing the object as the computed path plan uses the backprojected object 3D position to compute yaw angles for the robot. After detecting the object in the picture, we obtain a segmentation mask from detected boundary box using the Segment Anything model, and compute the center pixel of this mask. We feed this center pixel to the Boston dynamics grasping API to compute a motion plan to grasp the object.
Release Skill: We implement a simple routine for the release skill which takes two parameters: the object to be placed and the place receptacle. Once a navigation skill gets the robot to the place receptacle, the robot gently moves its arm up or down to release the grasped object, based on the place receptable 3D position. Future work will implement more complex semantic placement strategies to better leverage LIMP’s awareness and spatial grounding of instruction specific place receptacles. Kindly, visit our website to see demonstrations of these skills.
Appendix A7 Evaluation and Baseline Details
All computation including planning, loading and running pretrained visual language models was done on a single computer equipped with one NVIDIA GeForce RTX 3090 GPU. We leverage GPT-4-0613 as the underlying LLM for our instruction understanding module and all our baselines. In all experiments we set the LLM temperature to 0, however since deterministic greedy token decoding is not guaranteed with GPT4, we perform multiple (3) queries for each instruction and evaluate on the most recurring response (mode response).
We compare LIMP with baseline implementations of NLMap-Saycan [10] and Code-as-policies [12]. Both baselines use the same GPT-4 LLM, prompting structure, and in-context learning examples as our language understanding module. We integrate our composible referent descriptor syntax, spatial grounding module and low-level robot control into these baselines as APIs. This enables baselines to execute plans by querying relevant object positions, using our FMT* path planner to find paths to said positions and executing manipulation options.
A7.1 NLMap-Saycan Implementation Prompt
1You are an LLM for robot planning that understands logical operators such as &, |, ~ , etc. You have a finite set of robot predicates and spatial predicates, given a language instruction, your goal is to generate a sequence of actions that uses appropriate composition of robot and spatial predicates with relevant details from the instruction as arguments.
2Robot predicate set (near,pick,release).
3Usage:
4near[referent_1]:returns true if the desired spatial relationship is for robot to be near referent_1.
5pick[referent_1]:can only execute picking skill on referent_1 and return True when near[referent_1].
6release[referent_1,referent_2]:can only execute release skill on referent_1 and return True when near[referent_2].
7Spatial predicate set (isbetween,isabove,isbelow,isleftof,isrightof,isnextto,isinfrontof,isbehind).
8Usage:
9referent_1::isbetween(referent_2,referent_3):returns true if referent_1 is between referent_2 and referent_3.
10referent_1::isabove(referent_2):returns True if referent_1 is above referent_2.
11referent_1::isbelow(referent_2):returns True if referent_1 is below referent_2.
12referent_1::isleftof(referent_2):returns True if referent_1 is left of referent_2.
13referent_1::isrightof(referent_2):returns True if referent_1 is right of referent_2.
14referent_1::isnextto(referent_2):returns True if referent_1 is close to referent_2.
15referent_1::isinfrontof(referent_2):returns True if referent_1 is in front of referent_2.
16referent_1::isbehind(referent_2):returns True if referent_1 is behind referent_2.
17Rules:
18Strictly only use the finite set of robot and spatial predicates!
19Strictly stick to the usage format!
20Compose spatial predicates where necessary!
21You should strictly stick to mentioned objects, however you are allowed to propose and include plausible objects if and only if not mentioned in instruction but required based on context of instruction!
22Pay attention to instructions that require performing certain actions multiple times in generating and sequencing the predicates for the final Output!
23Example:
24Input_instruction: Go to the orange building but before that pass by the coffee shop, then go to the parking sign.
25Output:
261. near[coffee_shop]
272. near[orange_building]
283. near[parking_sign]
29Input_instruction: Go to the blue sofa then the laptop, after that bring me the brown bag between the television and the kettle on the left of the green seat, I am standing by the sink.
30Output:
311. near[blue_sofa]
322. near[laptop]
333. near[brown_bag::isbetween(television,kettle::isleftof(green_seat))]
344. pick[brown_bag::isbetween(television,kettle::isleftof(green_seat))]
355. near[sink]
366. release[brown_bag,sink]
37Input_instruction: Hey need you to pass by chair between the sofa and bag, pick up the bag and go to the orange napkin on the right of the sofa.
38Output:
391. near[chair::isbetween(sofa,bag)]
402. near[bag]
413. pick[bag]
424. near[orange_napkin::isrightof(sofa)]
43Input_instruction: Go to the chair between the green laptop and the yellow box underneath the play toy
44Output:
451. near[chair::isbetween(green_laptop,yellow_box::isbelow(play_toy))]
46Input_instruction: Check the table behind the fridge and bring two beers to the couch one after the other
47Output:
481. near[table::isbehind(fridge)]
492. pick[beer]
503. near[couch]
514. release[beer,couch]
525. near[table::isbehind(fridge)]
536. pick[beer]
547. near[couch]
558. release[beer,couch]
56Input_instruction: <given_instruction>
57Output:
Listing 5: Exact prompt to implement NlMap-Saycan LLM planner
A7.2 Code-as-Policies Implementation Prompt
1##Python robot planning script
2from robotactions import near, pick, release
3spatial_relationships = [
4"isbetween", #referent_1::isbetween(referent_2,referent_3):returns true if referent_1 is between referent_2 and referent_3.
5"isabove", #referent_1::isabove(referent_2):returns True if referent_1 is above referent_2.
6"isbelow", #referent_1::isbelow(referent_2):returns True if referent_1 is below referent_2.
7"isleftof", #referent_1::isleftof(referent_2):returns True if referent_1 is left of referent_2.
8"isrightof", #referent_1::isrightof(referent_2):returns True if referent_1 is right of referent_2.
9"isnextto", #referent_1::isnextto(referent_2):returns True if referent_1 is close to referent_2.
10"isinfrontof", #referent_1::isinfrontof(referent_2):returns True if referent_1 is in front of referent_2.
11"isbehind" #referent_1::isbehind(referent_2):returns True if referent_1 is behind referent_2.]
12##Rules:
13##Strictly only use the finite set of robot and spatial predicates!
14##Strictly stick to the usage format!
15##Compose spatial predicates where necessary!
16##You should strictly stick to mentioned objects, however you are allowed to propose and include plausible objects if and only if not mentioned in instruction but required based on context of instruction!
17##Pay attention to instructions that require performing certain actions multiple times in generating and sequencing the predicates for the final Output!
18# Go to the orange building but before that pass by the coffee shop, then go to the parking sign.
19ordered_navigation_goal_referents = ["coffee_shop", "orange_building", "parking_sign"]
20for referent in ordered_navigation_goal_referents:
21 near(referent)
22# Go to the blue sofa then the laptop, after that bring me the brown bag between the television and the kettle on the left of the green seat, I am standing by the sink.
23ordered_navigation_goal_referents = ["blue_sofa", "laptop", "brown_bag::isbetween(television,kettle::isleftof(green_seat))", "sink"]
24referents_to_pick= ["brown_bag::isbetween(television,kettle::isleftof(green_seat))"]
25release_location_referents = ["sink"]
26picked_item = None
27for referent in ordered_navigation_goal_referents:
28 near(referent)
29 if referent in referents_to_pick:
30 pick(referent)
31 picked_item = referent
32 if referent in release_location_referents:
33 release(picked_item, referent)
34#Hey need you to pass by chair between the sofa and bag, pick up the bag and go to the orange napkin on the right of the sofa.
35ordered_navigation_goal_referents = ["chair::isbetween(sofa,bag)", "bag", "orange_napkin::isrightof(sofa)"]
36referents_to_pick= ["bag"]
37picked_item = None
38for referent in ordered_navigation_goal_referents:
39 near(referent)
40 if referent in referents_to_pick:
41 pick(referent)
42 picked_item = referent
43#Go to the chair between the green laptop and the yellow box underneath the play toy
44near("chair::isbetween(green_laptop,yellow_box::isbelow(play_toy))")
45#Check the table behind the fridge and bring two beers to the couch one after the other
46for i in range(2):
47 near("table::isbehind(fridge)")
48 pick("beer")
49 near("couch")
50 release("beer", "couch")
51
52#<given_instruction>
Listing 6: Exact prompt to implement Code-as-policies planner
A7.3 Instruction set
We perform a large-scale evaluation on 150 instructions across five real-world environments. The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [16], and an additional 39 tasks with descriptive spatial constraints and temporal structures (CST).
11: put the red can in the trash bin
22: put the brown multigrain chip bag in the woven basket
33: find the succulent plant
44: pick up the up side down mug
55: put the apple on the macbook with yellow stickers
66: use the dyson vacuum cleaner"
77: bring the kosher salt to the kitchen counter
88: put the used towels in washing machine
99: move the used mug to the dish washer
1010: place the pickled cucumbers on the shelf
$1$: find my mug with the shape of a donut
1212: put the almonds in the almond jar
1313: fill the zisha tea pot with a coke from the cabinet
1414: take the slippery floor sign with you
1515: take the slippers that have holes on them to the shoe rack
1616: find the mug on the mini fridge
1717: bring the mint flavor gum to the small table
1818: find some n95 masks
1919: grab the banana with most black spots
2020: fill the empty bottle with lemon juice
2121: throw away the apple thats about to rot
$2$: throw away the rotting banana
2323: take the box of organic blueberries out of the fridge
2424: give a can of diet coke to the toy cat
Listing 7: Nlmap Detailed Object Tasks (NLMD)
11: I opened a pepsi earlier, bring an open can to the orange table
22: I spilled my coke, can you put a replacement on the kitchen counter
33: I spilled some coke on the television, go and bring something to clean it up
44: I accidentally dropped that jalapeno chips after eating it. Would you mind throwing it away
55: I like fruits, can you put something I would like on the yellow sofa for me
66: There is a green counter, a yellow counter, and a table. visit all the locations
77: There is a green counter, a trash can, and a table. visit all the locations
88: Redbull is my faviorite drink, can you put one on the desk please
99: Would you bring a coke can to the door for me
1010: Please, move the pepsi to the red counter
$1$: Can you move the coke can to the orange counter
1212: Would you throw away the bag of chips for me
1313: Put an energy bar and water bottle on the table
1414: Bring a lime soda and a bag of chips to the sofa
1515: Can you throw away the apple and bring a coke to the bed
1616: Bring a 7up can and a tea to the office desk
1717: Move the multigrain chips to the table and an apple to the yellow counter
1818: Move the lime soda, the sponge, and the water bottle to the table
1919: Bring two sodas to the table
2020: Move three cokes to the trash can
2121: Throw away two cokes from the counter
$2$: Bring two different sodas to the cabinet, there is a coke, pepsi, soup, tea and 7up in the fridge
2323: Bring an apple, a coke, and water bottle to the sofa
2424: I spilled my coke on the table, throw it away and then bring something to help clean
2525: I just worked out, can you bring me a drink and a snack to recover, i am on the sof
Listing 8: Nlmap Complex Language Understanding Tasks (NLMC)
11: Move the Takis on the desk to the nightstand
22: Move the soda can to the box
33: Move the purple shampoo to the red bag
44: Move the white meds box to the trash bin
55: Move the power adapter to the chair
66: Move the blue gloves to the sink
77: Move the McDonalds paper bag to the stove
88: Move the herbal tea can to the box
99: Move the cooking oil bottle to the marble surface
1010: Move the milk bottle to the chair
$1$: Move the purple shampoo to the white rack
1212: Move the purple lightbulb box to the sofa chair
1313: Pick up purple medicine, drop it on chair
1414: Pick up eyeglass case, drop it on chair
1515: Pick up grey rag, drop it in sink
1616: Pick up golden can rag, drop it on table
1717: Pick up red navaratna oil, drop it on table
1818: Pick up purple shampoo, drop it on green rack
1919: Pick up taki chips, drop it on countertop
2020: Pick up bandage box, drop it in dustbin
2121: Pick up white aerosol, drop it in trash can
$2$: Pick up peanut butter, drop it on countertop
2323: Pick up blue gloves, drop it in sink
2424: Pick up brown box, drop it on chair
2525: Pick up axe body spray, drop it on shel
Listing 9: Ok-Robot Tasks (OKRB)
11: go to brown bookshelf, metal desk, wooden desk, kitchen counter, and the blue couch in any order
22: move to grey door, then bookrack, then go to the brown desk, then counter, then white desk
33: visit brown wooden desk but only after bookshelf
44: go from brown bookshelf to white metal desk and only visit each landmark one time
55: go to brown wooden desk exactly once and do not visit brown desk before bookshelf
66: go to brown desk only after visiting bookshelf, in addition go to brown desk only after visiting white desk
77: visit the blue IKEA couch, in addition never go to the big steel door
88: visit white kitchen counter then go to brown desk, in addition never visit white table
99: go to the grey door, and only then go to the bookshelf, in addition always avoid the table
1010: go to kitchen counter then wooden desk, in addition after going to counter, you must avoid white table
$1$: Go to bookshelf, alternatively go to metal desk
1212: Go to counter, alternatively go to metal desk
1313: Go to the counter, but never visit the counter
1414: do not go to the wooden desk until bookshelf, and do not go to bookshelf until wooden desk
1515: go to brown desk exactly once, in addition go to brown desk at least twice
1616: move to couch exactly twice, in addition pass by counter at most once
1717: navigate to the counter then the brown desk, in addition after going to the counter, you must avoid doorway
1818: visit counter at least six times
1919: either go to bookshelf then the brown desk, or go to couch
2020: navigate to the wooden door, the glass door and the table, kitchen counter, and the blue couch in any order
2121: go to the painting, then find the kitchen table, then front desk, then staircase
$2$: navigate to classroom but do not visit classroom before the white table
2323: only visit classroom once, and do not visit classroom until you visit elevator first
2424: Go to the staircase, front desk and the white table in that exact order. You are not permitted to revisit any of these locations
2525: go to the front desk then the yellow office door, in addition do not visit the glass door
2626: go to the stairs then the front desk, in addition avoid purple elevator
2727: move to elevator then front desk, in addition avoid staircase
2828: go to front desk then the cabinet, always avoid the elevator
2929: Go to elevator, alternatively go to staircase
3030: Visit the elevator exactly once, in addition visit the front desk on at least 2 separate occasions
3131: Go to the office, in addition avoid visiting the elevator and the classroom
3232: Visit the front desk, in addition you are not permitted to visit elevator and staircase
$3$: Visit the purple door elevator, then go to the front desk and then go to the kitchen table, in addition you can never go to the elevator once you have seen the front desk
3434: Visit the front desk then the sofa then the white table, in addition if you visit the sofa you must avoid the television after that
3535: Go to the glass door, but never visit the glass door
3636: do not go to the white table until classroom, and do not go to the classroom until white table
3737: find the office, in addition avoid visiting the front desk and the classroom and the table
Listing 10: Lang2LTL Complex temporal Tasks (CT)
11: Go to the red sofa but dont pass by any door then go to the computer, the sofa is behind the pillar and on the right side
22: Find the door with posters in front of the yellow pillar then go stand by the tree but avoid any sofa
33: There is a cabinet, a television and a tree in this room. I want you to go to the last item I mentioned then pass by the second item but dont go close to the first when you are going to thsecond item.",
44: Go to the trash bin, after that go to the large television. Actually can you go to the laptop on the table before doing all that?
55: Try to go to the tree without going near the trash bin.
66: Find the yellow trash bin and go to it, then go by the white table. But before doing the first thing visit where the fridge is then after the last thing, stop by the red sofa but avoid theooden door with posters",
77: Go to the television but avoid the tree.
88: Go to any trash bin but not the yellow one
99: Visit one of the televisions in this room then go to the fridge to the left of the cabinet
1010: I need you to first go to the table between the pillar and the door with posters, then go to the tree. You know what, ignore the last thing I asked you to go to, after the first thing go tthe computer next to the green box instead.
$1$: Go to the red sofa then the blue one next to the cabinet but first pass by the book shelf
1212: Go to water filter, dont pass by any computer to the left of the red sofa
1313: Visit these things in the following order the microwave, then the black chair infront of the yellow robot then the fridge, but before doing any of this go to the robot
1414: Go to the cabinet between the blue sofa and the brown box.
1515: I want you to go the cabinet but dont go near any black chair. Hold on, actually can you pass by the whiteboard before doing all that?
1616: I need you to go to the red sofa but dont go anywhere near the sink
1717: You can either go as close as you can to the sink or the plant pot but dont go near the yellow robot.
1818: Go to the green curtain after that the large television but try not to go near the red sofa
1919: Go to the sink then the red sofa, then the guitar. After that I want you to return to the first thing you visited.
2020: I need you to go to the sink then any computer on a table but before you start that can you first go to the bookshelf, when you are done with all that, return to brown box
2121: There are a couple of things in this room, a coffee machine, a robot, a computer and a plant pot. Visit the second item I mentioned then the first, after that visit the brown box, then go the items you havent visited yet.
$2$: Avoid any cabinet but go get the mug on the blue sofa and take it to red one
2323: Please get me my plant pot from between the sofa and the sink, bring it to the white cabinet
2424: Can you pick up my book from the left side of the television and bring it to me? I am sitting on the couch which is to the right of the door with posters. Make sure to never go near the trhbin while doing all this, it on the yellow table.
2525: Bring the green plush toy to the whiteboard infront of it, watchout for the robot infront of the toy
2626: Go get my yellow bag and bring it to the table between the yellow pillar and the wooden door with posters next to the whiteboard
2727: Find the bottle that is on the table to the left of the computer and bring it to the wardrobe that is next to the glass door
2828: Go to the cabinet between the blue sofa and the yellow robot or the cabinet to the left of the blue sofa.
2929: Go take the green toy that is next to the sofa under the poster and bring it to the bookshelf
3030: I have a white cabinet, a green toy, a bookshelf and a red chair around here somewhere. Take the second item I mentioned from between the first item and the third. Bring it the cabinet butvoid the last item at all costs.
3131: Hey I am standing by the whiteboard infront of the bookshelf, can you bring me the mug from the table to the left of the fridge?
3232: Never pass by a robot, but i need you to bring the bag on the table to the cabinet
$3$: Take a soda can to the fridge, you can find one on the table to the left of the blue sofa
3434: Hey, can you pick up my bag for me? It is under the table infront of the glass door, bring it to the cabinet. Actually you know what, forget what i asked for and bring the green toy instea its at the same place",
3535: Go to the sink and take the mug beside the coffee maker, drop it off at the red sofa
3636: Go to the sofa behind the chair next to the orange counter in front of the pillar
3737: I want you to go visit the chair behind the sofa then the green bag, do that 3 times but during the second time avoid the fridge
3838: Go bring the plush toy between the sofa and the television to the couch, my cat is on the brown table so please dont pass near the table when you are returning with the toy
3939: I think I left my wallet on the kitchen counter, go get it i will meet you at the bed. There is a lantern near the bed, make sure you dont hit it
Listing 11: Complex Spatiotemporal Tasks (CST)

### v3

@@
 Abstract
-Enabling mobile robots to follow complex natural language instructions is an important yet challenging problem.
-People want to flexibly express constraints, refer to arbitrary landmarks and verify behavior when instructing robots.
-Conversely, robots must disambiguate human instructions into specifications and ground instruction referents in the real world.
-We propose Language Instruction grounding for Motion Planning (LIMP), an approach that enables robots to verifiably follow expressive and complex open-ended instructions in real-world environments without prebuilt semantic maps.
-LIMP constructs a symbolic instruction representation that reveals the robot’s alignment with an instructor’s intended motives and affords the synthesis of robot behaviors that are correct-by-construction.
-We perform a large scale evaluation and demonstrate our approach on 150 instructions in five real-world environments showing the generality of our approach and the ease of deployment in novel unstructured domains.
-In our experiments, LIMP performs comparably with state-of-the-art LLM task planners and LLM code-writing planners on standard open vocabulary tasks and additionally achieves 79% success rate on complex spatiotemporal instructions while LLM and Code-writing planners both achieve 38%.
-See supplementary materials and demo videos at robotlimp.github.io.
-Keywords: Foundation Models, Mobile Manipulation, Task and Motion Planning
-1 Introduction
-Robots need a rich understanding of natural language to be usable by non-experts in unstructured environments.
+When instructing robots, users want to flexibly express constraints, refer to arbitrary landmarks, and verify robot behavior, while robots must disambiguate instructions into specifications and ground instruction referents in the real world.
+To address this problem, we propose Language Instruction grounding for Motion Planning (LIMP), an approach that enables robots to verifiably follow complex, open-ended instructions in real-world environments without prebuilt semantic maps.
+LIMP constructs a symbolic instruction representation that reveals the robot’s alignment with an instructor’s intended motives and affords the synthesis of correct-by-construction robot behaviors.
+We conduct a large-scale evaluation of LIMP on 150 instructions across five real-world environments, demonstrating its versatility and ease of deployment in diverse, unstructured domains.
+LIMP performs comparably to state-of-the-art baselines on standard open-vocabulary tasks and additionally achieves a 79% success rate on complex spatiotemporal instructions, significantly outperforming baselines that only reach 38%.
+[^1] See supplementary materials and demo videos at robotlimp.github.io
+I Introduction
+[…]
+Robots need a rich understanding of natural language to be instructable by non-experts in unstructured environments.
 People, on the other hand, need to be able to verify that a robot has understood a given instruction and will act appropriately.
-However, natural language instructions can be quite complex, often featuring ambiguous phrasing, intricate spatiotemporal constraints, and unique referents.
-Consider the instruction: “Go to the building with windows while avoiding the red picnic tables, then bring me the book between the pine tree and the graffiti wall”, to solve this task a robot must ground open-vocabulary referents (“building with windows”, “red picnic tables”, “book”, “pine tree”, “graffiti wall”), adhere to temporal constraints (“then bring”, “while avoiding”), and disambiguate multiple instances of an object category using descriptive spatial information (“book between the pine tree and the graffiti wall”).
-The rise of foundation models in robotics [1, 2] offers a path to achieving such long-horizon goals, and indeed some approaches for robot instruction following have been proposed, but have largely been focused on navigation [3, 4, 5, 6, 7].
-These methods, broadly classified under object goal navigation [8], enable navigation to instances of an object category, but are limited in localizing spatial references and disambiguating object instances based on descriptive language.
-Other works [9, 10, 11] that extend instruction following to mobile manipulation have been limited to tasks with simple temporal constraints expressed in unambiguous language.
-Existing efforts typically rely on Large Language Models (LLMs) as complete planners, bypassing intermediate symbolic representations that may provide verification of correctness prior to task execution.
-Alternative approaches that leverage code-writing LLMs [5, 6, 12] are susceptible to errors in generated code which may lead to unsafe/unaligned robot behaviors.
-Borrowing ideas from formal methods to map between natural language and specification languages like temporal logics [13] provides a robust framework for language disambiguation, handling complex temporal constraints, and behavior verification.
+Achieving these objectives, however, is challenging as natural language instructions often feature ambiguous phrasing, intricate spatiotemporal constraints, and unique referents.
+To illustrate, consider the instruction shown in Figure 1: “Bring the green plush toy to the whiteboard in front of it, watch out for the robot in front of the toy”.
+Solving such a task requires a robot to ground open-vocabulary referents, follow temporal constraints, and disambiguate objects using spatial descriptions.
+Foundation models [1, 2] offer a path to achieving such complex long-horizon goals; however, existing approaches for robot instruction following have largely focused on navigation [3, 4, 5, 6, 7].
+These methods, broadly classified under object goal navigation [8], enable navigation to instances of an object category but are limited in their ability to localize spatial references and disambiguate object instances based on descriptive language.
+Other works [9, 10, 11] extend instruction following to mobile manipulation but are limited to tasks with simple temporal constraints expressed in unambiguous language.
+Moreover, existing efforts typically rely on Large Language Models (LLMs) as complete planners, bypassing intermediate symbolic representations that could provide verification of correctness before execution.
+Alternative approaches leveraging code-writing LLMs [5, 6, 12] are susceptible to errors in generated code, which may lead to unsafe robot behaviors.
+Mapping natural language to specification languages like temporal logic [13] provides a robust framework for language disambiguation, handling complex temporal constraints, and behavior verification.
 However, prior works along this line require prebuilt semantic maps with discrete sets of prespecified referents/landmarks from which instructions can be constructed [7, 14, 15].
+Figure 1: Our approach executing the instruction “Bring the green plush toy to the whiteboard in front of it, watch out for the robot in front of the toy”.
+The robot dynamically detects and grounds open-vocabulary referents with spatial constraints to construct an instruction-specific semantic map, then synthesizes a task and motion plan to solve the task.
+In this example, the robot navigates from its start location (yellow, A), to the green plush toy (green, B), executes a pick skill then navigates to the whiteboard (blue, C), and executes a place skill.
+Note that the robot has no prior semantic knowledge of the environment.
 We propose Language Instruction grounding for Motion Planning (LIMP), a method that leverages foundation models and temporal logics to dynamically generate instruction-conditioned semantic maps that enable robots to construct verifiable controllers for following navigation and mobile manipulation instructions with open vocabulary referents and complex spatiotemporal constraints.
-In a novel environment, LIMP first constructs a 3D representation of the environment via SLAM, then leverages large language models to translate complex natural language instructions into temporal logic specifications with a novel composable syntax that enables referent disambiguation.
-Instruction referents are then detected and grounded using Visual Language Models (VLMs) and spatial reasoning.
-Finally, a task and motion plan is synthesized to progressively guide the robot through the subgoals required to satisfy the instruction, as shown in Figure LABEL:fig:splash.
-In summary, we make the following contributions: (1) A modular framework for translating expressive natural language instructions into temporal logic, grounding instruction referents, and executing the command via task and motion planning (TAMP).
+In a novel environment, LIMP constructs a 3D map via SLAM, then uses LLMs to translate complex natural language instructions into temporal logic specifications with a novel composable syntax for referent disambiguation.
+Instruction referents are detected and grounded using vision-language models (VLMs) and spatial reasoning.
+Finally, a task and motion plan is synthesized to guide the robot through the required subgoals, as shown in Figure 1.
+In summary, we make the following contributions: (1) A modular framework that translates expressive natural language instructions into temporal logic, grounds instruction referents, and executes commands via Task and Motion Planning (TAMP).
 (2) A spatial grounding method for detecting and localizing open vocabulary objects with spatial constraints in 3D metric maps.
-(3) A TAMP algorithm that localizes regions of interest (goal/avoidance zones) and synthesizes constraint-satisfying motion plans to progressively solve long-horizon tasks.
-2 Related Works
-In this section, we briefly highlight the most relevant comparable works in the space of visual scene understanding, instruction following from natural language, and task and motion planning.
-We offer a more detailed literature review in our appendix A2.
-Visual Scene Understanding: NLMap [10] presents a scene representation that grounds open-vocabulary language queries to spatial locations via pre-trained VLMs.
-While NLMap is suitable for handling complex descriptions of individual objects, it is fundamentally unable to handle instructions involving complex constraints between multiple objects since it does not capture object relationships.
-LIMP handles these more complicated instructions by using a novel spatial grounding module to incorporate spatial relationships between objects, and leverages task and motion planners to satisfy these constraints.
-Natural Language Instruction Following: Lang2LTL [16] is a multi-stage LLM-based approach that performs entity extraction and replacement to translate natural language instructions into a temporal logic expression.
-However, Lang2LTL relies on a prebuilt semantic map that grounds expression symbols, whereas our approach uses a novel scene representation that supports open-vocabulary language and generates the relevant landmarks based on the open-vocabulary instruction.
-Task and Motion Planning: Action-Oriented Semantic Maps (AOSMs) [17] are a class of semantic maps that include additional models in space for where the robot can execute manipulation skills (represented as symbols), and are designed to work in conjunction with TAMP solvers for mobile manipulation domains.
-Our scene representation is similar to an AOSM since it captures spatial information about semantic objects of interest, and is compatible with TAMP solvers.
-However, once an AOSM is generated for a scene, there is only a closed-set of goals that can be planned for, whereas our approach can handle open-vocabulary task specifications.
-3 Background
+(3) A TAMP algorithm that localizes regions of interest (goal/avoidance zones) and synthesizes constraint-satisfying motion plans for long-horizon tasks.
+II Background and Related Works
+We briefly highlight the most relevant works in visual scene understanding [10], natural language instruction following [7, 16], and task and motion planning [17], and provide a comprehensive review in our supplementary materials.
+NLMap [10] grounds open-vocabulary language queries to spatial locations using pre-trained VLMs.
+While effective for describing individual objects, it cannot handle instructions involving complex constraints between multiple objects due to the lack of object relationship modeling.
+LIMP addresses this with a novel spatial grounding module that resolves spatial relationships and leverages task and motion planners to satisfy these constraints.
+Lang2LTL [7] is a multi-stage, LLM-based approach that uses entity extraction and replacement to translate language instructions into temporal logic.
+Its extension [16] incorporates VLMs and semantic information (via text embeddings) to ground referents.
+These works require prebuilt semantic maps/databases describing landmarks to ground symbols, whereas our approach dynamically generates landmarks based on open-vocabulary instructions.
+Action-Oriented Semantic Maps (AOSMs) [17] augment semantic maps with models indicating where robots can perform manipulation skills, integrating with TAMP solvers for mobile manipulation.
+LIMP similarly provides a TAMP-compatible spatial representation but supports open-vocabulary tasks, whereas AOSMs remain constrained to a fixed set of goals once generated.
+II-A Linear Temporal Logic
 LIMP translates natural language instructions into temporal logic specifications for verifiable task and motion planning.
-Our approach is compatible with various specification languages and planning frameworks, however we choose Linear Temporal Logic (LTL) [18] to represent instructions due to its proven expressivity in representing a wide range of complex robot mission requirements [19].
-LTL offers a rich and expressive syntax for defining the temporal properties of systems.
-Specifications consist of atomic propositions that reflect an aspect of a system’s state and the logical operators: negation ($\neg$), conjunction ($\land$), disjunction ($\lor$), implication ($\rightarrow$), in addition to temporal operators: next ($\mathcal{X}$), until ($\mathcal{U}$), globally ($\mathcal{G}$) and finally ($\mathcal{F}$).
-Despite its expressivity, LTL has seen limited use due to the expert knowledge required to construct specifications, however recent works have seen significant success removing this barrier by directly translating natural language into LTL [7, 20, 14, 21, 22, 23].
+While compatible with various specification languages and planning frameworks, we choose Linear Temporal Logic (LTL) [18] for its proven expressivity in representing complex robot mission requirements [19].
+LTL defines temporal properties using atomic propositions, logical operators—negation ($\neg$), conjunction ($\land$), disjunction ($\lor$), implication ($\rightarrow$)—and temporal operators: next ($\mathcal{X}$), until ($\mathcal{U}$), globally ($\mathcal{G}$), and finally ($\mathcal{F}$).
+Despite its expressivity, LTL has been underutilized due to the expert knowledge required to construct specifications, however recent works have seen significant success directly translating natural language into LTL [7, 20, 14, 21, 22, 23].
 Behavior Verification: Expressing instructions as temporal logic specifications allows us to verify the correctness of generated plans a priori.
-However, instead of secondary explicit verification via approaches such as model checking, we follow conventional wisdom [24] and directly use specifications to synthesize plans that are correct-by-construction [25, 26].
-4 Problem Definition
-Given a natural language instruction $l$, our goal is to synthesize and sequence navigation and manipulation behaviors to produce a policy that satisfies the temporal and spatial constraints expressed in $l$.
-Spatial constraints determine the success of the task based on the sequence of poses the robot traversed during task execution, and temporal constraints determine the sequencing of spatial constraints as a function of task progression.
-We assume that a robot equipped with an RGBD camera has navigated around some space and captured a collection of images along with camera poses, with this data we build a metric map $m$ (e.g: pointcloud, 3D voxelgrid) of the environment.
-This map defines the space of possible $SE(3)$ poses $P$ and can be used to localize the robot (i.e: estimate $p_{robot}\in P$).
-Noteably, unlike previous work that leverage temporal logics [16], we do not assume access to a semantic map with the locations of objects or predicates prespecified.
-Instead, we leverage two types of foundation models: a task-agnostic visual language model $\sigma$ that given an image and text provides a bounding box/segmentation of the image based on the text, and an auto-regressive large language model $\psi$ that samples likely language tokens based on a history of tokens.
-Navigation: We formalize navigation as object-goal oriented continuous path planning problems, where the goal is to generate paths to a goal pose set $P_{goals}\subset P$ while staying in feasible regions ($P_{feasible}\subset P$) and avoiding infeasible regions ($P_{infeasible}=P_{feasible}^{C}$).
+However, instead of explicit verification methods such as model checking, we leverage insights from prior works [24] and directly use specifications to synthesize plans that are correct-by-construction [25, 26].
+III Problem Definition
+Given a natural language instruction $l$, our goal is to synthesize and sequence navigation and manipulation behaviors to produce a policy that satisfies the temporal and spatial constraints in $l$.
+Spatial constraints determine task success based on the sequence of robot poses traversed during execution; temporal constraints determine the sequencing of these spatial constraints as a function of task progression.
+We assume a robot with an RGB-D camera has already navigated a space, capturing images and camera poses.
+From this data, we build a metric map $m$ (e.g., point cloud, 3D voxel grid) of the environment, defining the space of possible $SE(3)$ poses $P$ and enabling robot localization (i.e., estimating $p_{\text{robot}}\in P$).
+Unlike previous work leveraging temporal logic [7], we do not assume access to a semantic map with prespecified object locations or predicates.
+Instead, we leverage two foundation models: a task-agnostic vision-language model $\sigma$ that, given an image and text, provides bounding boxes or segmentations based on the text; and an auto-regressive large language model $\psi$ that samples likely language tokens based on a history of tokens.
+Navigation: Navigation is formalized as an object-goal oriented continuous path planning problem, where the goal is to generate paths to a goal pose set $P_{goals}\subset P$ while staying in feasible regions ($P_{feasible}\subset P$) and avoiding infeasible regions ($P_{infeasible}=P_{feasible}^{C}$).
 Infeasible regions include environment obstacles as well as dynamically determined semantic regions that violate constraints in the instruction $l$.
-Manipulation: We formalize interactive manipulation behaviours as options [27] parameterized by objects.
+Manipulation: We formalize manipulation behaviours as options [27] parameterized by objects.
 Consider an object parameter $\theta$ that parameterizes an option $O_{\theta}=(I_{\theta},\pi_{\theta},\beta_{\theta})$, the initiation set, policy, and termination condition are functions of both the robot pose $P$ and $\theta$.
 The initiation set $I_{\theta}$ denotes the global reference frame robot positions and object-centric attributes––such as object size––that determine if the option policy $\pi_{\theta}$ can be executed on the object $\theta$.
-To execute a manipulation skill on an object, an object-goal navigation behavior must first be synthesized and executed to bring the robot into proximity with the object in question.
-Our formalism is capable of encoding a wide variety of long-horizon object-centric tasks, but for simplicity, we demonstrate our approach with two types of tasks: multi-object goal navigation and open vocabulary mobile pick-and-place [9, 28].
-These tasks require navigation, picking, and placing robot behaviors.
-5 Language Instruction Grounding for Motion Planning
+To execute a manipulation skill on an object, an object-goal navigation behavior must first be executed to bring the robot into proximity with the object.
+We assume access to a library of these manipulation skills and demonstrate our approach on multi-object goal navigation and open-vocabulary mobile pick-and-place [9, 28].
+IV Language Instruction Grounding for Motion Planning
 Figure 2: [A] LIMP translates natural language instructions into temporal logic expressions, where open-vocabulary referents are applied to predicates that correspond to robot skills––note the context-aware resolution of the phrase “blue one” to the referent “blue_sofa”.
-[B] Visual language models detect referents, while spatial reasoning disambiguates referent instances to generate a 3D semantic map that localizes instruction-specific referents.
-[C] Finally, the temporal logic expression is compiled into a finite-state machine, which a bi-level task and motion planner uses with dynamically-generated task progression semantic maps to progressively identify goals and constraints in the environment, and generate a plan that satisfies the high-level task specification.
-LIMP interprets expressive natural language instructions and generates instruction-conditioned semantic maps that enable robots to verifiably solve long-horizon tasks with complex spatiotemporal constraints (Figure 1).
-Our approach has a modular structure consisting of a Language Instruction Module, a Spatial Grounding Module and a Task and Motion Planning Module.
-The first two modules utilize pretrained foundation models, while the third involves a task and motion planner for sequencing task subgoals and generating constraint-satisfying motion plans.
-5.1 Language Instruction Module
-Figure 2: An instruction is first translated into a conventional LTL formula $\phi_{l}$ that loosely captures the desired temporal occurrence of referent objects, then into our LTL syntax $\varphi_{l}$ with predicate functions that temporally chain required robot skills parameterized by composible referent descriptors.
-In this module, we leverage a large language model $\psi$ to translate a natural language instruction $l$ into a linear temporal logic specification $\varphi_{l}$ with a novel composable syntax that enables referent disambiguation.
-We achieve this with a two-stage in-context learning strategy, the first stage translates $l$ into a conventional LTL formula $\phi_{l}$ where propositions refer to open-vocabulary objects.
-The second stage then takes as input $l$ and $\phi_{l}$ to generate a new formula $\varphi_{l}$, with predicate functions that correspond to parameterized robot skills.
-We define three predicate functions: near, pick and release for the primitive navigation and manipulation skills (picking and placing) required for multi-object goal navigation and mobile pick-and-place.
-Predicate functions in $\varphi_{l}$ are parameterized by Composable Referent Descriptors (CRDs), our novel propositional expressions that represent specific referent instances by chaining comparators that encode descriptive spatial information.
-For example the statement “the yellow cabinet above the fridge that is next to the stove”, can be represented with the composible referent descriptor below:
+[B] Vision-language models detect referents, while spatial reasoning disambiguates referent instances to generate a 3D semantic map that localizes instruction-specific referents.
+[C] Finally, the temporal logic expression is compiled into a finite-state automaton, which a task and motion planner uses with dynamically-generated task progression semantic maps to progressively identify goals and constraints in the environment, and generate a plan that satisfies the high-level task specification.
+LIMP interprets expressive natural language instructions to generate instruction-conditioned semantic maps, enabling robots to verifiably solve long-horizon tasks with complex spatiotemporal constraints (Figure 2).
+We briefly describe our modular approach in this section and present comprehensive implementations details in our supplementary materials.
+IV-A Language Instruction Module
+In this module, we leverage a large language model $\psi$ to translate a natural language instruction $l$ into a linear temporal logic specification $\varphi_{l}$ with a novel composable syntax for referent disambiguation.
+We achieve this through a two-stage in-context learning strategy.
+The first stage prompts $\psi$ to translate $l$ into a conventional LTL formula $\phi_{l}$ where propositions refer to open-vocabulary objects.
+The second stage takes $l$ and $\phi_{l}$ as input and prompts $\psi$ to generate a new formula $\varphi_{l}$ with predicate functions corresponding to parameterized robot skills.
+We define three predicate functions—near, pick, and release—for the primitive navigation and manipulation skills required for multi-object goal navigation and mobile pick-and-place.
+Predicate functions in $\varphi_{l}$ are parameterized by Composable Referent Descriptors (CRDs), our novel propositional expressions representing specific referent instances by chaining comparators that encode descriptive spatial information.
+For example, the instruction “the yellow cabinet above the fridge that is next to the stove” can be represented with the CRD:
 […]
-which specifies that there is a fridge that is next to a stove, and the specific yellow cabinet we desire is above that fridge.
+This specifies that there is a fridge next to a stove, and the desired yellow cabinet is above that fridge.
 CRDs are constructed from a set of 3D spatial comparators [29] defined in our prompting strategy.
-It is important to note that unlike recent works [9, 30], instructions do not require any specific phrasing syntax nor specific keywords and may feature arbitrary complexity and ambiguity.
-The LLM $\psi$ directly samples the entire LTL formula $\varphi_{l}$ with predicate functions parameterized by composable referent descriptors with appropriate spatial comparators based on the context of the instruction.
-Figure 2 presents an example highlighting the result of our two-stage prompting strategy.
-See appendix section A3 for further details.
-LLM Verification: Verifying the LTL formula $\varphi_{l}$ sampled from the LLM is vital as errors in referent extraction and temporal task structure implied by the sampled formula adversely affect accurate instruction following.
-To that end, the symbol verification node highlighted in Figure 1 leverages the unique properties of LTL to provide high-level human-in-the-loop verification for the extracted instruction referents and temporal task structure.
-Recent work [31] has addressed providing ISO 61508 [32] standard safety guarantees in robot task execution by translating descriptions of safety constraints from natural language to LTL formulas which are verified to be correct by human experts and used to enforce robot behavior.
-Similarly, we rely on human checking to verify that the translated formula $\varphi_{l}$ is correct, the symbol verification node implements an interactive dialog system that presents users with the extracted referent CRDs in the LTL formula as well as the implied task structure and reprompts the LLM based on corrective statements of the user to obtain new formulas.
-Unlike [31] we improve readability and relax the need for experts by directly translating the task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task––back into English statements via a simple deterministic translation scheme.
-In our experiments (Table 1 & 2), we find that even without this human verification and reprompting step, the initial formulas sampled by our language understanding module impressively encodes the correct referents and implied temporal task structure.
-See appendix section A3.1 for further details.
-5.2 Spatial Grounding Module
+Unlike recent works [9, 11], our approach does not require specific phrasing or keywords and can handle instructions with arbitrary complexity and ambiguity.
+The LLM $\psi$ directly samples the entire LTL formula $\varphi_{l}$ with predicate functions parameterized by CRDs using appropriate spatial comparators based on the instruction’s context.
+Figure 3 illustrates the result of our two-stage prompting strategy.
+Figure 3: An instruction is first translated into a conventional LTL formula $\phi_{l}$ that loosely captures the desired temporal occurrence of referent objects, then into our LTL syntax $\varphi_{l}$ with predicate functions that temporally chain required robot skills parameterized by composable referent descriptors.
+LLM Verification: Verifying the LTL formula $\varphi_{l}$ sampled from the LLM is crucial as errors in referent extraction and temporal task structure affects instruction following accuracy.
+Our symbol verification node (Figure 2) leverages LTL properties to provide high-level human-in-the-loop verification of extracted instruction referents and temporal task structure.
+Recent work [30] provides ISO 61508 [31] safety guarantees in robot task execution by translating safety constraints from natural language to LTL formulas, which are verified by human experts and used to enforce robot behavior.
+Similarly, we rely on human verification to ensure the translated formula $\varphi_{l}$ is correct.
+Our symbol verification node implements an interactive dialog system that presents users with the extracted referent CRDs and implied task structure, and reprompts the LLM based on user corrections to obtain new formulas.
+Unlike prior work [30], we eliminate the need for experts by directly translating the task structure—encoded in the LTL formulas’s equivalent automaton—back into English statements via a simple deterministic translation scheme.
+In our experiments (Tables I and II), we find that even without human verification and reprompting, the initial formulas sampled by our language understanding module impressively encode the correct referents and temporal task structure.
+IV-B Spatial Grounding Module
 Figure 4: [A] Our spatial grounding module leverages a VLM to detect all referent occurrences from prior observations of the environment.
-[B] We then obtain an initial semantic map with all detected referent instances by backprojecting pixels in segmented referent masks unto the 3D map.
-[C] We resolve each referent’s spatial comparators with respect to the origin coordinate frame of reference.
-[D] We finally filter out failing instances to obtain a Referent Semantic Map (RSM) that localizes the exact referent instances described in the instruction.
-The spatial grounding module detects and localizes specific instances of objects referenced in a given instruction.
-To achieve this we extract composable referent descriptors from translated LTL formulae and leverage visual language models [33, 34] to detect and generate segmentation masks of all referent occurrences from the robot’s prior observations of the environment.
-We then backproject pixels in these masks unto our 3D map to obtain an initial semantic map that localizes all instruction object instances.
-In the example from Figure 2, all occurrences of green_plush_toy, whiteboard and robot are detected, segmented and backprojected unto the map as shown in Figure 3[a&b].
-We filter detected occurrences of referents to obtain the specific object instances described in the given instruction by resolving the 3D spatial comparators captured in each referent’s CRD, recall that CRDs are propositional expressions and can be evaluated as true or false.
-We define eight (8) spatial comparators (isbetween, isabove, isbelow, isleftof, isrightof, isnextto, isinfrontof, isbehind) to reason about the spatial relationships between referents based on their backprojected 3D positions.
+[B] An initial semantic map with all detected referent instances is generated by backprojecting pixels in segmented referent masks unto the 3D map.
+[C] Each referent’s spatial comparators is resolved with respect to the origin coordinate frame of reference.
+[D] Failing instances are filtered out to obtain a Referent Semantic Map (RSM) that localizes the exact referent instances described in the instruction.
+This module detects and localizes specific object instances referenced in an instruction.
+From the translated LTL formulas, we extract composable referent descriptors (CRDs) and use vision-language models OWL-ViT [32] and SAM [33] to detect and segment all referent occurrences from the robot’s prior observations of the environment.
+We backproject pixels in these segmentation masks onto our 3D map, creating an initial semantic map of all instruction object instances.
+From the example in Figure 3, occurrences of green_plush_toy, whiteboard, and robot are detected, segmented, and backprojected onto the map (Figure 4[a&b]).
+To obtain the specific object instances described in the instruction, we resolve the 3D spatial comparators in each referent’s CRD––recall that CRDs are propositional expressions and can be evaluated as true or false.
+We define eight spatial comparators (isbetween, isabove, isbelow, isleftof, isrightof, isnextto, isinfrontof, isbehind) to reason about spatial relationships based on backprojected 3D positions.
 Since all backprojected positions are relative to an origin coordinate system, our spatial comparators are resolved from the perspective of this origin position as shown in Figure 4[c].
-This type of Relative frame of reference (FoR) when describing spatial relationships between objects, in contrast to an Absolute or Intrinsic FoR, is dominant in English [35], hence a logical choice in our work.
-For each detected referent, the backprojected 3D position of its center mask pixel is used as its representative position and we resolve a given referent with a spatial description by applying the appropriate spatial comparator to all detected pairs of the desired referent and comparison landmark objects.
-This filtering process yields a Referent Semantic Map (RSM) that localizes specific object instances described in the original instruction as shown in Figure 3[d].
-See appendix section A4 for further details.
+This type of relative frame of reference (FoR) when describing spatial relationships between objects, in contrast to an absolute or intrinsic FoR, is dominant in English [34], and is a logical choice for our work.
+Using the 3D position of each referent’s center mask pixel as its representative position, we resolve a given referent with a spatial description by applying the appropriate spatial comparator to all detected pairs of the desired referent and comparison landmark objects.
+This filtering process yields a Referent Semantic Map (RSM) that localizes specific object instances described in the instruction as shown in Figure 4[d].
 VLM Verification: Potential misclassifications from object detector VLMs is the main source of error in this module.
-We do not address interactively correcting VLM misclassifications as that is out of the scope of this work, however we provide 3D visualizations tools that enable end users to visually inspect and verify that constructed referent semantic maps localized referents correctly.
-5.3 Task and Motion Planning Module
+We do not address interactively correcting VLM misclassifications as that is out of the scope of this work, but we provide 3D visualization tools that enable users to visually inspect and verify that constructed referent semantic maps correctly localize referents.
+IV-C Task and Motion Planning Module
 Figure 5: [A] A given instruction translated into our LTL syntax $\varphi_{l}$ can be compiled into an equivalent finite-state automaton that captures the temporal constraints of the task.
-We select a path through this automaton with a strategy that incrementally selects the next progression state from the initial state to the accepting state.
-We then execute manipulation options and navigation behaviors as required by this high-level task plan.
-[B] To execute navigation objectives we generate a task progression semantic map (TPSM) that augments the environment with state transition constraints, localizing goal (yellow) and avoidance (red) regions.
-We convert generated TPSMs into 2D obstacle maps for constraint-aware continuous path planning.
+A path through this automaton is selected with a strategy that incrementally picks the next progression state from the initial state to the accepting state.
+The robot then executes the manipulation options and navigation behaviors dictated by this high-level task plan.
+[B] To execute navigation objectives our approach generates a task progression semantic map (TPSM) that augments the environment with state transition constraints, localizing goal (yellow) and avoidance (red) regions.
+Generated TPSMs are converted into 2D obstacle maps for constraint-aware continuous path planning.
 Finally, our TAMP module synthesizes and sequences navigation and manipulation behaviors to produce a plan that satisfies the temporal and spatial constraints expressed in the given instruction.
-Progressive Motion Planner: Our TAMP algorithm (Alg.1) first compiles an LTL formula with parameterized robot skills into an equivalent finite-state automaton (Figure 4[a]), which can be used to generate a plan that is guaranteed to be correct-by-construction.
-We select a path that traverses this automaton from its initial state to its accepting state, this is a high-level task plan that interleaves the navigation and manipulation objectives required for satisfying the instruction.
-We implement a simple strategy that incrementally selects the next progression state from the initial state until the final accepting state is reached, ensuring the generated plan faithfully obeys all temporal subgoal objectives.
-As shown in Figure 4[a] each progressive automaton state is connected by transition edges that represent the logical expression that must hold for the transition to occur.
-Our algorithm performs the required low-level behavior to achieve each transition, for manipulation objective subgoals, we execute the appropriate parameterized skill.
-For navigation objective subgoals, we dynamically generate Task Progression Semantic Maps (TPSM) that localize goal and constraint regions in the environment and perform continuous path planning using the Fast Marching Tree algorithm (FMT) [36].
-Task Progression Semantic Maps: A TPSM (Alg.2) augments a 3D scene with the navigation constraints specified by logical state transition expressions, enabling a robot to perform goal-directed and constraint-aware navigation.
-Segmented regions of interest captured in a TPSM can be arbitrarily large via a nearness threshold (in meters) that defines what it means to be at or close to an object.
-This parameter can be specified as a high-level parameter or included in the language instruction module’s prompting strategy, such that an LLM infers what value to use based on the given instruction.
-Goal regions captured in a generated TPSM along with object-centric constraints, such as object shape and size, form the initiation set for manipulation options.
-Similar to our spatial grounding module, TPSMs are agnostic to temporal logic task representations and can be used out of the box with a variety of planning approaches to localise regions of interest and enable semantic constraint aware motion planning.
-We primarily evaluate our approach on a ground mobile robot, hence we transform 3D TPSMs into 2D geometric obstacle maps, where constraint regions are also treated as obstacle regions (Figure 4[b]).
-Our approach however, is robot agnostic and allows for direct planning in 3D TPSMs for appropriate embodiments, such as drones.
-6 Experimental Evaluation
-Table 1: Performance comparison of one-shot instruction understanding and spatial resolution.
+Progressive Motion Planner (PMP): Our TAMP algorithm compiles the LTL formula with parameterized robot skills into an equivalent finite-state automaton (Figure 5[a]) to generate a verifiably correct task and motion plan.
+A path from the initial to the accepting state in this automaton is a high-level task plan that interleaves navigation and manipulation objectives required to satisfy the instruction.
+We select such a path with a simple strategy that incrementally selects the next progression state until the accepting state is reached, ensuring the plan obeys all temporal subgoal objectives.
+As shown in Figure 5[a], automaton states are connected by transition edges representing the logical expressions required for transitions.
+For each transition, our algorithm executes the necessary low-level behaviors: for manipulation subgoals, it executes the appropriate parameterized skill; for navigation subgoals, it dynamically generates Task Progression Semantic Maps (TPSMs) to localize goal and constraint regions and performs continuous path planning using the Fast Marching Tree algorithm (FMT[35].
+Task Progression Semantic Maps (TPSM): A TPSM augments a 3D scene with navigation constraints specified by logical state transition expressions, enabling goal-directed, constraint-aware navigation.
+Regions of interest in a TPSM are defined using a nearness threshold specifying proximity to an object.
+This threshold can be set globally or included in the language instruction module’s prompting strategy, allowing an LLM to infer its value based on the instruction.
+Like our spatial grounding module, TPSMs are agnostic to temporal logic representations and can be used with various planning approaches for semantic constraint-aware motion planning.
+We primarily evaluate our approach on a ground mobile robot, hence we transform 3D TPSMs into 2D geometric obstacle maps, where constraint regions are treated as obstacles (Figure 5[b]).
+However, our approach is robot-agnostic and supports direct planning in 3D TPSMs for appropriate embodiments like drones.
+V Evaluation
+Our evaluations test the hypothesis that translating natural language instructions into LTL expressions and dynamically generating semantic maps enables robots to accurately interpret and execute instructions in large-scale environments without prior training.
+We focus on three key questions: (1) Can our language instruction module interpret complex, ambiguous instructions?
+(2) Can our spatial grounding module resolve specific object instances described in instructions?
+(3) Can our TAMP algorithm generate constraint-satisfying plans?
+To answer these questions, we compare LIMP with two baselines: an LLM task planner (NLMap-Saycan [10]) and an LLM code-writing planner (Code-as-Policies [12]), representing state-of-the-art approaches for language-conditioned, open-ended robot task execution.
+Both baselines use the same LLM (GPT-4-0613), prompting structure, and in-context learning examples as our language instruction module.
+In Code-as-Policies, in-context examples are converted into language model-generated program (LMP) snippets [12].
+To ensure competitive performance, we integrate our CRD syntax, spatial grounding module, and low-level robot control into these baselines, allowing them to query object positions, use our FMT path planner, and execute manipulation skills.
+We also evaluate ablations of our two-stage language instruction module due to its importance in instruction following.
+In our full approach, the first stage prompts an LLM to generate a conventional LTL formula $\phi_{l}$ from instruction $l$ by dynamically selecting relevant in-context examples from a standard dataset [14] based on cosine similarity.
+Our first ablation selects in-context examples randomly; and the second ablation skips this stage entirely, directly sampling our LTL syntax with parameterized robot skills $\varphi_{l}$ from $l$.
+We conduct a large-scale evaluation across five real-world environments on a diverse task set of 150 instructions from multiple prior works [10, 11, 7].
+This task set consists of 24 tasks with fine-grained object descriptions (NLMD), 25 tasks with complex language (NLMC), 25 tasks with simple structured phrasing (OKRB), 37 tasks with complex temporal structures (CT) and 39 tasks with descriptive spatial constraints and temporal structures (CST).
+Below are examples from each task category illustrating the variety in complexity:
+1NLMD: Put the brown multigrain chip bag in the woven basket
+2NLMC: I like fruits, can you put something I would like on the yellow sofa for me
+3OKRB: Move the soda can to the box
+4CT: Visit the purple door elevator, then go to the front desk and then go to the kitchen table, in addition you can never go to the elevator once you have seen the front desk
+5CST: I have a white cabinet, a green toy, a bookshelf and a red chair around here somewhere.
+Take the second item I mentioned from between the first item and the third.
+Bring it the cabinet but avoid the last item at all costs.
+To evaluate instruction understanding, we introduce performance metrics: referent resolution accuracy, avoidance constraint resolution accuracy, and spatial relationship resolution accuracy.
+These metrics utilize the word error rate (WER), widely used in speech recognition to quantify the difference between a reference and a hypothesis transcription by computing the minimal number of substitutions, deletions, and insertions needed to transform the hypothesis into the reference.
+WER is calculated as $WER=\frac{S+D+I}{N}$, where $S$ is substitutions, $D$ deletions, $I$ insertions, and $N$ is the total number of words in the reference.
+In our work:
+  • Referent resolution accuracy compares extracted referents in the generated LTL formula to ground truth referents.
+  • Avoidance constraint resolution accuracy compares referents to avoid in the LTL formula (denoted by the unary negation operator) against ground truth avoidance referents.
+  • Spatial relationship resolution accuracy compares generated Composable Referent Descriptors (CRDs) in the LTL specification with ground truth CRD expressions.
+We also define temporal alignment accuracy and planning success rate.
+A plan is temporally aligned if the sequence of subgoals matches the instructor’s intention, and successful if it satisfies all spatial and temporal constraints specified in the instruction.
+Achieving a high plan success rate is challenging, requiring accurate referent and avoidance constraint resolution, spatial grounding, and temporal alignment.
+We report average word error rates for each baseline in Table I and the percentage of successful and temporally accurate plans in Table II.
+TABLE I: Performance comparison of one-shot instruction understanding and spatial resolution.
 Approach Referent Resolution Accuracy (Average WER) $\downarrow$ Avoidance Constraint Resolution Accuracy (Average WER) $\downarrow$ Spatial Relationship Resolution Accuracy (Average WER) $\downarrow$
@@
 Limp Two Stage Prompting [Similar Embedding] 0.07 0.04 0.03
-Table 2: Performance comparison of one-shot temporal alignment and plan success rate.
+TABLE II: Performance comparison of one-shot temporal alignment and plan success rate.
 Approach Temporal Alignment Accuracy (of instructions) $\uparrow$ Planning Success Rate (of instructions) $\uparrow$
@@
 Limp Two Stage Prompting [Similar Embedding] 88% 80% 100% 76% 92% 79% 76% 100% 65% 79%
-Our evaluations test the hypothesis that our system for translating natural language instructions into LTL expressions and dynamically generating semantic maps enables robots to accurately interpret and execute instructions in large-scale environments without prior training in said environments.
-We focus on three key questions: (1) Can our language module interpret complex ambiguous natural language instructions?
-(2) Can our spatial grounding module resolve specific object instances described in instructions?
-(3) Can our TAMP algorithm generate constraint-satisfying plans?
-Figure 5: Our interactive reprompting strategy implemented in the symbol verification node regenerates corrective formulas that improve plan success rates with minimal chat turns.
-To answer these questions, we compare LIMP with baseline implementations of an LLM task planner (NLMap-Saycan) [10] and an LLM code-writing planner (Code-as-policies) [12], representing state-of-the-art approaches for language-conditioned open ended robot task execution.
-Both baselines use the same underlying LLM (gpt-4-0613), prompting structure, and in-context learning examples as our language understanding module.
-For the code-as-policies planner, in-context examples are converted into language model-generated program (LMP) snippets [12].
-To ensure competitive performance, we integrate our CRD syntax, spatial grounding module and low-level robot control into these baselines, allowing them to query object positions, use our FMT* path planner, and execute manipulation options.
-We also evaluate ablations of our two-stage language understanding module, due to the prominence of language understanding on instruction following.
-In our complete approach, the first stage samples a conventional LTL formula $\phi_{l}$ from an instruction $l$ by dynamically selecting relevant in-context examples from a standard dataset [14] of natural language and LTL pairs based on cosine similarity.
-Our first ablation selects in-context examples randomly, and our second ablation skips this stage entirely, directly sampling our LTL syntax with parameterized robot skills $\varphi_{l}$.
-Figure 6: [A] Sample generated plan for a multi object-goal navigation task. [B] Sample generated plan for a mobile pick-and-place task.
-We perform a large-scale evaluation on 150 instructions across five real-world environments.
-The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [16], and an additional 39 tasks with descriptive spatial constraints and temporal structures (CST).
-To evaluate instruction understanding, we introduce the following performance metrics: Referent resolution accuracy, Avoidance constraint resolution accuracy, and Spatial relationship resolution accuracy.
-These metrics use the popular word error rate metric (WER) to measure alignment with the ground-truth.
-Additionally, we define Temporal alignment accuracy and Planning success rate metrics.
-A plan is temporally aligned if the sequence of subgoal objectives matches the instructor’s intention, and it is deemed successful if it satisfies all spatial and temporal constraints specified in the instruction.
-Achieving a high plan success rate requires accurate referent and avoidance constraint resolution, spatial grounding, and temporal alignment.
-We report the average word error rates for each baseline in Table 1 and the percentage of successful plans and temporally accurate plans in Table 2.
-In addition to the verification benefits of symbolic planning, LIMP outperforms baselines in most tasksets, especially in complex temporal planning and constraint avoidance.
-We observe that LLM and code-writing planners generate sequential subgoals effectively, but struggle with temporal constraint adherence.
-In contrast, our approach ensures each robot step adheres to constraints while achieving subgoals.
-Additional qualitative results are shown in Figure 6, and robot demonstrations are available on our project website.
-7 Limitations and Conclusion
-Our study revealed several limitations that merit discussion.
-LIMP is able to interpret non-finite instructions into LTL formulas however our planner is only able to processes co-safe formulas, which are limited to finite sequences.
-The accuracy of spatial grounding is dependent on VLM object recognition and depth sensing, with any deficiencies affecting performance.
-LIMP assumes a consistent environment between mapping and execution, making our approach non-reactive to dynamic changes.
-Addressing this with editable scene representations is a future goal.
-Moreover, while our Progressive Motion Planning algorithm is complete, it does not provide any optimality guarantees, future work will focus on enhancing planning efficiency.
-Foundation models have the potential to drive the next generation of generally capable robots.
-Our results suggest leveraging these models for their strengths (LLMs for language, VLMs for vision) in combination with proven approaches for safety, explainability, and verifiable behavior synthesis.
-This approach for using foundation models in robotics is vital to improve safety and trust in robots that will have to work with and for humans.
-Supplementary Materials
-Appendix A1 Appendix Summary
+VI Discussion
+Beyond the verification benefits of symbolic planning, LIMP outperforms baselines in most task sets, notably in complex temporal planning and constraint avoidance.
+While NLMap-Saycan and Code-as-Policies effectively generate sequential subgoals, they struggle with strict temporal constraints—for example, avoiding a specific referent while approaching another.
+Our approach ensures each robot step adheres to constraints while achieving subgoals, explaining LIMP’s superior performance on CT and CST tasks.
+As shown in Table II, LIMP underperforms only against NLMap-Saycan in the NLMC task category.
+This task set, introduced in the same paper as the baseline [10] (which outperforms LIMP), includes instructions with implicit details such as: “I like fruits, can you put something I would like on the yellow sofa for me.” NLMap-Saycan is better suited to infer and generate plans with possible fruit options, whereas our few-shot LTL translation process is not designed for this.
+VII Limitations and Conclusion
+Although LIMP is capable of interpreting non-finite instructions into LTL formulas, our planner is currently limited to processing co-safe formulas, which handle only finite sequences.
+The accuracy of spatial grounding relies on the performance of vision-language models (VLMs) for object recognition meaning any shortcomings in these systems can negatively impact results.
+Additionally, LIMP assumes a static environment between mapping and execution, making it not responsive to dynamic changes—an area we aim to address with future work on editable scene representations.
+Our Progressive Motion Planning algorithm is complete but does not guarantee optimality; however, our framework can be used with existing TAMP planners to enhance efficiency.
+Foundation models hold significant promise for advancing the next generation of autonomous robots.
+Our results suggest that combining these models—LLMs for language and VLMs for vision—with established methods for safety, explainability, and verifiable behavior synthesis can lead to more reliable and capable robotic systems.
+Acknowledgement
+This work was supported by the Office of Naval Research (ONR) under REPRISM MURI N000142412603 and ONR #N00014-22-1-2592, as well as the National Science Foundation (NSF) via grant #1955361.
+Partial funding was also provided by The Robotics and AI Institute.
+Appendix
+A1 Appendix Summary
 These sections presents additional details on our approach for leveraging foundation models and temporal logics to verifiably follow expressive natural language instructions with complex spatiotemporal constraints without prebuilt semantic maps.
 We encourage readers to visit our website robotlimp.github.io for project summary and demonstration videos.
-Appendix A2 Extended Related Works
-A2.1 Foundation Models in Robotics
+A2 Extended Related Works
+A2-A Foundation Models in Robotics
 Grounding language referents to entities and actions in the world [36, 37, 38, 39] is challenging in part due to the fact that complex perceptual and behavioral meaning can be constructed from the composition of a wide-range of open-vocabulary components[40, 20, 21, 41, 37].
@@
 Additionally, the symbols in our temporal logic translation correspond to parameterized task relevant robot skills as opposed to propositions of referent entities extracted from instructions.
-A2.2 Planning Models in Robotics
+A2-B Planning Models in Robotics
 Semantic Maps: Semantic maps [49] are a class of scene representations that capture semantic (and typically geometric) information about the world, and can be used in cojunction with planners to generate certain types of complex robot behavior like collision-free navigation with spatial constraints [50, 51].
@@
 LIMP contains sufficient information to produce a problem and domain description augmented with geometric information for bi-level TAMP solvers like [54, 55].
-Appendix A3 Language Instruction Module
+A3 Language Instruction Module
 We implement a two-stage prompting strategy in our language instruction module to translate natural language instructions into LTL specifications.
@@
 Listing 2: Second stage prompt to output our LTL syntax with CRD parameterized robot skills
-A3.1 Interactive Symbol Verification
+A3-A Interactive Symbol Verification
 Verifying sampled LTL formulas is essential, as such we implement an interactive dialog system that presents users with extracted referent composible referent descriptors (CRDs) in sampled formulas as well as the implied task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task.
@@
 Users can verify a formula as correct or provide corrective statements which are used to reprompt the LLM to obtain new formulas.
-Here is the exact text prompt used for reprompting.
+Below is the exact prompt used for regenerating formulas.
 1There was a mistake with your output LTL formula: Error with <verification_type>!
@@
 Listing 3: Corrective reprompting prompt used to obtain new LTL formulas
-As an illustration, the instruction “Bring the green plush toy to the whiteboard in front of it” yields the interactive Referent and Task Structure Verification dialog shown in Figure A.7.
-Figure A.7: Sample interactive referent and task structure verification dialog
-Appendix A4 Spatial Grounding Module
+To illustrate, the instruction “Bring the green plush toy to the whiteboard in front of it” yields the interactive Referent and Task Structure Verification dialog below:
+1**************************
+2Instruction Following
+3**************************
+4Input_instruction: "Bring the green plush toy to the whiteboard in front of it"
+5Sampled LTL formula: F(A & F(B & F(C & FD)))
+6 A: near[green_plush_toy]
+7 B: pick[green_plush_toy]
+8 C: near[whiteboard::isinfrontof(green_plush_toy)]
+9 D: release[green_plush_toy, whiteboard::isinfrontof(green_plush_toy)]
+10
+11***************************
+12Referent Verification
+13***************************
+14I extracted this list of relevant objects based on your instruction:
+15 * whiteboard::isinfrontof(green_plush_toy)
+16 * green_plush_toy
+17Does this match your intention? (y/n)
+18
+19****************************
+20Task Structure Verification
+21****************************
+22Based on my understanding here is the sequence of subgoal objectives needed to satisfy the task:
+23Subgoal_1:
+24 Logical Expression: A&!B
+25 English translation: I should be near the [green_plush_toy] and not have picked up the [green_plush_toy]
+26Subgoal_2:
+27 Logical Expression: B&!C
+28 English translation: I should have picked up the [green_plush_toy] and not be near the [whiteboard::isinfrontof(green_plush_toy)]
+29Subgoal_3:
+30 Logical Expression: C&!D
+31 English translation: I should be near the [whiteboard::isinfrontof(green_plush_toy)] and not have released the [green_plush_toy] at the [whiteboard::isinfrontof(green_plush_toy)]
+32Subgoal_4:
+33 Logical Expression: D
+34 English translation: I should have released the [green_plush_toy] at the [whiteboard::isinfrontof(green_plush_toy)]
+35Does this match your intention? (y/n)
+Listing 4: Interactive referent and task structure verification dialog.
+A4 Spatial Grounding Module
 The spatial grounding module detects and localizes specific instances of objects referenced in a given instruction by first detecting, segmenting and back-projecting all referent occurances and then filtering based on the descriptive spatial details captured by each referent’s composable referent descriptor (CRD).
@@
 $S=\{w\in W\mid\exists g\in G\text{ such that }\text{isinfrontof}(w,g)\text{ istrue}\}$ (A.3)
-A4.1 3D Spatial Comparators
+A4-A 3D Spatial Comparators
 Our 3D spatial comparators enable Relative Frame of Reference (FoR) spatial reasoning between referents, based on their backprojected 3D positions.
@@
 88. isbehind(referent_1_pos, referent_2_pos, threshold): Returns true if the x-coordinate of referent_1 exceeds that of referent_2 by at least threshold’, indicating referent_1 is behind referent_2.
-Listing 4: Implementation description of 3D spatial comparators
-Appendix A5 Task and Motion Planning Module
+Listing 5: Implementation description of 3D spatial comparators
+A5 Task and Motion Planning Module
 We present pseudocode for our Progressive Motion Planner (Alg.1) and our algorithm for generating Task Progression Semantic Maps (Alg.2).
@@
 The points are then integrated into $\mathcal{M}_{\text{tpsm}}$, yielding a semantic map that identifies goal and constraint violating regions.
-A5.1 Algorithm Pseudocode
 Algorithm 1 Progressive Motion Planning Algorithm
@@
 32:end procedure
-Appendix A6 Robot Skills
+A6 Robot Skills
 We define three predicate functions: near, pick and release for the navigation, picking and placing skills required for multi-object goal navigation and mobile pick-and-place.
@@
 We discuss navigation at length in the paper, so here we focus on the pick and place manipulation skills.
-Pick Skill: Once the robot has executed the near skill and is at the object to be manipulated, we take a picture of the current environment to detect the object using the Owl-Vit model.
+Pick Skill: Once the robot has executed the near skill and is at the object to be manipulated, it takes a photo of the current environment to detect the object using the Owl-Vit model.
 The robot is guaranteed to be facing the object as the computed path plan uses the backprojected object 3D position to compute yaw angles for the robot.
@@
 Kindly, visit our website to see demonstrations of these skills.
-Appendix A7 Evaluation and Baseline Details
+A7 Evaluation and Baseline Details
 All computation including planning, loading and running pretrained visual language models was done on a single computer equipped with one NVIDIA GeForce RTX 3090 GPU.
 We leverage GPT-4-0613 as the underlying LLM for our instruction understanding module and all our baselines.
-In all experiments we set the LLM temperature to 0, however since deterministic greedy token decoding is not guaranteed with GPT4, we perform multiple (3) queries for each instruction and evaluate on the most recurring response (mode response).
+In all experiments we set the LLM temperature to 0, however since deterministic greedy token decoding is not guaranteed with GPT4, we perform three (3) queries for each instruction and evaluate on the most recurring response (mode response).
 We compare LIMP with baseline implementations of NLMap-Saycan [10] and Code-as-policies [12].
@@
 This enables baselines to execute plans by querying relevant object positions, using our FMT* path planner to find paths to said positions and executing manipulation options.
-A7.1 NLMap-Saycan Implementation Prompt
+We visualize some qualitative results of LIMP from our experiments in Figure A.1.
+We also highlight results in Figure A.2 that illustrates how our interactive symbol verification and reprompting strategy A3-A improves instruction satisfaction with minimal chat turns for different instruction sets.
+Figure A.1: [A] Sample generated plan for a multi object-goal navigation task. [B] Sample generated plan for a mobile pick-and-place task.
+Figure A.2: Our interactive reprompting strategy implemented in the symbol verification node regenerates corrective formulas that improve plan success rates with minimal chat turns.
+A7-A NLMap-Saycan Implementation Prompt
 1You are an LLM for robot planning that understands logical operators such as &, |, ~ , etc.
@@
 57Output:
-Listing 5: Exact prompt to implement NlMap-Saycan LLM planner
-A7.2 Code-as-Policies Implementation Prompt
+Listing 6: Exact prompt to implement NlMap-Saycan LLM planner
+A7-B Code-as-Policies Implementation Prompt
 1##Python robot planning script
@@
 52#<given_instruction>
-Listing 6: Exact prompt to implement Code-as-policies planner
-A7.3 Instruction set
+Listing 7: Exact prompt to implement Code-as-policies planner
+A7-C Instruction set
 We perform a large-scale evaluation on 150 instructions across five real-world environments.
-The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [16], and an additional 39 tasks with descriptive spatial constraints and temporal structures (CST).
+The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [7], and an additional 39 tasks we propose that have descriptive spatial constraints and temporal structures (CST).
 11: put the red can in the trash bin
@@
 2424: give a can of diet coke to the toy cat
-Listing 7: Nlmap Detailed Object Tasks (NLMD)
+Listing 8: Nlmap Detailed Object Tasks (NLMD)
 11: I opened a pepsi earlier, bring an open can to the orange table
@@
 2525: I just worked out, can you bring me a drink and a snack to recover, i am on the sof
-Listing 8: Nlmap Complex Language Understanding Tasks (NLMC)
+Listing 9: Nlmap Complex Language Understanding Tasks (NLMC)
 11: Move the Takis on the desk to the nightstand
@@
 2525: Pick up axe body spray, drop it on shel
-Listing 9: Ok-Robot Tasks (OKRB)
+Listing 10: Ok-Robot Tasks (OKRB)
 11: go to brown bookshelf, metal desk, wooden desk, kitchen counter, and the blue couch in any order
@@
 3737: find the office, in addition avoid visiting the front desk and the classroom and the table
-Listing 10: Lang2LTL Complex temporal Tasks (CT)
+Listing 11: Lang2LTL Complex temporal Tasks (CT)
 11: Go to the red sofa but dont pass by any door then go to the computer, the sofa is behind the pillar and on the right side
@@
 3939: I think I left my wallet on the kitchen counter, go get it i will meet you at the bed. There is a lantern near the bed, make sure you dont hit it
-Listing 11: Complex Spatiotemporal Tasks (CST)
+Listing 12: Complex Spatiotemporal Tasks (CST)